### v1

# Variational Learning of Disentangled Representations

Abstract
Disentangled representations enable models to separate factors of variation that are shared across experimental conditions from those that are condition-specific. This separation is essential in domains such as biomedical data analysis, where generalization to new treatments, patients, or species depends on isolating stable biological signals from context-dependent effects. While extensions of the variational autoencoder (VAE) framework have been proposed to address this problem, they frequently suffer from leakage between latent representations, limiting their ability to generalize to unseen conditions. Here, we introduce DISCoVeR, a new variational framework that explicitly separates condition-invariant and condition-specific factors. DISCoVeR integrates three key components: (i) a dual-latent architecture that models shared and specific factors separately; (ii) two parallel reconstructions that ensure both representations remain informative; and (iii) a novel max-min objective that encourages clean separation without relying on handcrafted priors, while making only minimal assumptions. Theoretically, we show that this objective maximizes data likelihood while promoting disentanglement, and that it admits a unique equilibrium. Empirically, we demonstrate that DISCoVeR achieves improved disentanglement on synthetic datasets, natural images, and single-cell RNA-seq data. Together, these results establish DISCoVeR as a principled approach for learning disentangled representations in multi-condition settings.
1 Introduction
Neural network-based models have shown impressive success in learning informative representations of complex data, and are increasingly used in scenarios where each data point $x\in\mathcal{X}\subseteq\mathbb{R}^{d}$ is associated with a condition label $y\in{1,\dots,K}$. In biological applications, for example, the condition may correspond to a treatment, patient, or species. In biological contexts, for instance, the condition might reflect a treatment, a patient, or a species. The learned representations often support downstream tasks such as domain adaptation and transfer learning (e.g., Pan et al. (2010)), enabling models to generalize from observed conditions to new, related ones. This requires learning disentangled data representations that separate components of $x$ shared across all conditions from those specific to a given condition $y$.
In recent years, generative models have emerged as a powerful class of methods for representation learning, thanks to their ability to model complex data distributions and capture underlying latent structure. Prominent examples include Generative Adversarial Networks (GANs) (Goodfellow et al., 2020), Variational Autoencoders (VAEs) (Kingma and Welling, 2014), and diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020). Among these, VAEs and their extensions have become particularly prominent in applications involving transfer learning and domain adaptation (Akrami et al., 2020; Lovrić et al., 2021; Godinez et al., 2022; Zhang et al., 2023), due to their principled probabilistic formulation and their ability to represent uncertainty in learned features.
Several recent VAE-based approaches have been proposed to integrate data from multiple conditions or sources (Xu et al., 2021; Lotfollahi et al., 2019; Boyeau et al., 2022). However, only few explicitly aim to disentangle shared from condition-specific components (Sohn et al., 2015; Klys et al., 2018; Joy et al., 2020). While these methods offer notable improvements in disentanglement, they either depend on carefully crafted priors, difficult to specify in high-dimensional complex settings like single-cell genomics, or fail to achieve a clean separation, limiting their utility in applications that demand robust generalization across conditions.
In this work, we propose a new framework for learning disentangled representations in multi-condition datasets. Our contributions are as follows: (i) We propose a new method based on two distinct reconstruction objectives combined with adversarial learning. Unlike prior approaches, our framework minimizes reliance on restrictive priors or handcrafted components, relying less restrictive assumptions. (ii) We formulate the task of learning disentangled representations as max-min optimization problem, derive a corresponding objective and give theoretical guarantees for the equilibrium of this objective. (iii) Through experiments on synthetic benchmarks and real-world datasets in computer vision and single-cell biology, we show that our approach consistently achieves better disentanglement of shared and condition-specific structure compared to existing methods.
2 DISCoVeR: DISentangling COmmon and Variant Representations
For the task of learning disentangled representations from multi-condition data, we consider a dataset $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{n}$ consisting of inputs $x_{i}\in\mathcal{X}\subseteq\mathbb{R}^{d}$ collected from associated condition labels $y_{i}\in{1,\dots,K}$. We assume that for each class $k$, corresponding to a study or experimental condition, and that the associated subset $\mathcal{D}_{k}\coloneqq\{x_{i}:y_{i}=k\}$ consists of i.i.d. samples drawn from a class-conditional distribution $p(x|y=k)$.
2.1 Model assumptions
We assume that the data is generated by latent variables $z$ and $w$, such that the joint distribution $p(x,y,z,w)$ factorizes according to the probabilistic graphical model illustrated in Figure 1(a), i.e.,
$p(x,y,z,w)=p(y)p(w\mid y)p(z)p(x\mid z,w).$ (1)
$y$$w$$z$$x$
$y$$\,w\mid x,y\,$$\,z\mid x\,$$\,\hat{x}\mid z\,$$\hat{y}$$x$$\,\tilde{x}\mid z,w\,$
Figure 1: Graphical overview of our model. (a) Probabilistic graphical model: gray circles denote observed variables, white show latent variables. (b) Encoder–decoder architecture: the inhibition arrow from $\hat{y}$ to $\hat{x}$ corresponds to the adversarial component.
This model encodes two key conditional independence assumptions:
  1. 1.
    Latent variable conditional independence: Given the condition $y$, the latent representations $z$ and $w$ are conditionally independent: $z\perp w\mid y$.
  2. 2.
    Sufficiency of the shared latent representation: The input $x$ is conditionally independent of the condition $y$ given $w$: $x\perp y\mid w$.
However, note that in this model, $z$ and $w$ are no longer independent if conditioned also on $x$, that is, $z\not\perp w\mid x,y$.
2.2 Target posterior structure
In our model, the observed data $x$ is generated from two latent variables, $z$ and $w$. The marginal distribution of $z$ is independent of the condition label $y$, making $z$ condition-invariant, whereas $w$ depends on $y$ and is therefore condition-aware. Our goal is to learn probabilistic data representations such that the marginal distributions of $z$ and $w$ preserve the structure $z\perp y$ and $w\not\perp y$, and therefore yield disentangled representations. For this, we approximate the joint posterior $p_{z,w|x,y}$ via variational inference, enforcing that $z$ remains independent of $y$, while $w$ retains dependence on $y$. Our approach follows the empirical Bayes paradigm (Robbins, 1956), and in particular aligns with the principles of $f$-modeling introduced by Efron (2014), where the prior structure is inferred from the data to ensure consistency with observed marginals.
Ideally we would approximate the true posterior $p_{z,w|x,y}$ with a variational distribution $q_{z,w|x,y}$. However, even under simple variational families such as multivariate Gaussians, capturing dependencies between $z$ and $w$ requires modeling a full covariance structure, which becomes computationally intractable in high dimensions. To mitigate this, we approximate the posterior $p_{z,w|x,y}$ with a factorized variational distribution $q_{z|x}\ q_{w|x,y}$.
Our variational approximation is guided by two complementary objectives: (i) $q_{z|x}$ closely approximating the marginal posterior $p_{z|x}$; and (ii) the product $q_{z|x}\ q_{w|x,y}$ closely approximating the true posterior $p_{z,w|x,y}$. Formally, given variational families [^1] Here we consider general families and specify our concrete choices in Section 2.4. $\mathcal{Q}_{z}$ and $\mathcal{Q}_{w}$ we seek to find $q^{*}_{z|x}\in\mathcal{Q}_{z}$ and $q^{*}_{w|x,y}\in\mathcal{Q}_{w}$ that minimize the following sum of Kullback-Leibler (KL) divergences:
$q^{*}_{z|x},\ q^{*}_{w|x,y}=\arg\min_{\begin{subarray}{c}q_{z|x}\in\mathcal{Q}_{z}\\ q_{w|x,y}\in\mathcal{Q}_{w}\end{subarray}}\operatorname{D_{\mathrm{KL}}}(q_{z|x}\ \|\ p_{z|x})+\operatorname{D_{\mathrm{KL}}}(q_{z|x}q_{w|x,y}\ \|\ p_{z,w|x,y}).$ (2)
2.3 Optimization objective
Direct computation of the KL divergences is Equation 2 is intractable. Therefore, we introduce a surrogate objective composed of three terms, two of which are evidence lower bound (ELBO) terms that minimize the sum of KL divergences without requiring their direct evaluation, and an additional adversarial term.
The corresponding ELBO objective to minimize $\operatorname{D_{\mathrm{KL}}}(q_{z|x}\ \|\ p_{z|x})$ is
$\mathcal{L}_{z}(q_{z|x},p;x)\coloneqq\mathbb{E}_{q_{z|x}}\left[\log p(x\mid z)\right]-\operatorname{D_{\mathrm{KL}}}(q_{z|x}\,\|\,p_{z}),$ (3)
and the ELBO objective for the second KL term, $\operatorname{D_{\mathrm{KL}}}(q_{z|x}q_{w|x,y}\ \|\ p_{z,w|x,y})$, is given by
$\textstyle\mathcal{L}_{w}(q_{w|x,y},p;x,y)\!\coloneqq$ $\mathbb{E}_{q_{z|x}}\!\left[\mathbb{E}_{q_{w|x,y}}\!\left[\log\left(p(x|z,w)\right)\right]\right]\!\!-\!\operatorname{D_{\mathrm{KL}}}\left(q_{z|x}\,\|\,p_{z}\right)\!\!-\!\operatorname{D_{\mathrm{KL}}}\left(q_{w|x,y}\,\|\,p_{w\mid y}\right).$ (4)
Note that $\mathcal{L}_{w}(q_{w|x,y},p;x,y)$ is the ELBO objective that corresponds to a factorized posterior $q_{z|x}q_{w|x,y}$. In Proposition 2.1 we examine the gap between this objective, and an ELBO term corresponding to a full variational posterior $q_{z,w|x,y}$. This can be interpreted as the cost of enforcing a condition-invariant latent representation, specifically, constraining $z$ to depend only on $x$.
Proposition 2.1.
For random variables $x,y,z,w$ following the graphical model in Figure 1(a),
$\operatorname{ELBO}(q,p;x,y)-\mathcal{L}_{w}(q_{w|x,y},p;x,y)=\operatorname{D_{\mathrm{KL}}}\left(q_{z|x}\,\|\,p_{z|w,x,y}\right).$
where $\operatorname{ELBO}(q,p;x,y)\coloneqq\log p(x\mid y)-\operatorname{D_{\mathrm{KL}}}\left(q_{w|x,y}\,\|\,p_{w|x,y}\right)$.
The proof is provided in Appendix B.1.
Note that a full definition of the objectives in Equations 3 and 4 requires the specification of corresponding prior distributions, namely $p_{z}$ and $p_{w|y}$. We defer their definitions to Section 2.4.
The ELBO objective in Equation 4 is defined over the conditional log-likelihood $\log p(x\mid y)$, and requires the addition of $\log p(y)$ to correspond to the log-marginal likelihood $\log p(x,y)$. However, since $\log p(y)$ does not depend on the variational distributions, including it in the objective has no effect on optimization. Still, we wish to ensure that the implicit marginal over $y$—as inferred from the latent representations, matches the true marginal $p(y)$. To this end, we introduce an auxiliary classifier $g(y\mid z)$ as a form of posterior regularization (Ganchev et al., 2010). This classifier captures the residual predictive signal about $y$ contained in $z$, and is trained via the expected negative log-likelihood $-\mathbb{E}_{q(z\mid x)}\log g(y\mid z)$. If $z$ is indeed independent of $y$, then $g(y\mid z)$ will approximate the marginal distribution $p(y)$. By penalizing deviations from this behavior, we enforce the structural constraint $z\perp y$ in the learned representation.
Note that for this term to effectively encourage $q_{z|x}$ to discard condition-specific information, the classifier $g_{y|z}\in\mathcal{G}$ must be optimized separately from the the rest of the model. This prevents degenerate solutions in which the third term is minimized without truly removing information about $y$ from $z$, for example, by collapsing $g$ to always predict a constant output, regardless of $z$.
Combining the three terms above, we define the objective
$\mathcal{L}(q_{z|x},q_{w|x,y},g_{y|z};x,y)=\mathcal{L}_{z}(q_{z|x},p;x)+\mathcal{L}_{w}(q_{w|x,y},p;x,y)-\mathbb{E}_{q_{z|x}}\log g(y\mid z),$ (5)
which can be explicitly expressed as
$\mathcal{L}(q_{z|x},q_{w|x,y},g_{y|z};x,y)\coloneqq$ $\mathbb{E}_{q_{z|x}}[\log p(x\mid z)]+\mathbb{E}_{q_{z|x}}\left[\mathbb{E}_{q_{w|x,y}}[\log p(x\mid z,w)]\right]$
$-\mathbb{E}_{q_{z|x}}[\log g(y\mid z)]-2\operatorname{D_{\mathrm{KL}}}(q_{z|x}\,\|\,p_{z})-\operatorname{D_{\mathrm{KL}}}\left(q_{w|x,y}\,\|\,p_{w\mid y}\right).$ (6)
Finally, to enable flexible trade-offs between reconstruction expressiveness and disentanglement, we introduce weighting terms $\alpha=(\alpha_{1},\alpha_{2})$ into the objective following the motivation of $\beta$-VAEs (Higgins et al., 2017):
$\mathcal{L}_{\alpha}(q_{z|x},q_{w|x,y},g_{y|z};x,y)\coloneqq$ $\mathbb{E}_{q_{z|x}}[\log p(x\mid z)]+\mathbb{E}_{q_{z|x}}\left[\mathbb{E}_{q_{w|x,y}}[\log p(x\mid z,w)]\right]$ (7)
$-\mathbb{E}_{q_{z|x}}\log g(y\mid z)-\alpha_{1}\operatorname{D_{\mathrm{KL}}}(q_{z|x}\,\|\,p_{z})-\alpha_{2}\operatorname{D_{\mathrm{KL}}}\left(q_{w|x,y}\,\|\,p_{w\mid y}\right).$
Accordingly, the mean weighted objective is suitable for max-min optimization of the form:
$\max_{q_{z|x}\in\mathcal{Q}_{z}}\max_{q_{w|x,y}\in\mathcal{Q}_{w}}\min_{g_{y|z}\in\mathcal{G}}\mathbb{E}_{p_{x,y}}\left[\mathcal{L}_{\alpha}(q_{z|x},q_{w|x,y},g_{y|z};x,y)\right].$ (8)
2.4 Latent prior models and variational approximations
Prior specification
We place a standard normal prior over the condition-invariant latent variable, $p_{z}=\mathcal{N}(0,I)$, which reflects a non-informative prior belief over its values.
For the condition-aware latent variable $w$, we define a class-conditional Gaussian prior $p_{w|y}$. As a simple specification, we model $z$ and $w$ to be of the same dimension, and specify
$p(w\mid y=k)=\mathcal{N}(\mu_{k},I)$ (9)
$\mu_{k}\coloneqq\mathbb{E}_{p_{x|y=k}}\left[\mathbb{E}_{q_{z|x}}[z]\right].$ (10)
Here $\mu_{k}$ is the mean of the inferred latent representations $z$ within each class. Similarly, for different dimensions of $z$ and $w$ the mean aggregation can be replaced with a neural network that maps the inferred representations $z$ for each class to the parameters of the Gaussian prior.
This specification induces a coupling between the two latent variables through the data distribution. Aligning $p_{w|y}$ with the class-wise expectations of the invariant variable, further encourages $q_{z|x}$ to encode informative representations, since capturing the shared structures will now not only increase $\mathcal{L}_{z}(q_{z|x},p;x)$, but also decrease $\operatorname{D_{\mathrm{KL}}}\left(q_{w\mid x,y}\,\|\,p_{w|y}\right)$, and thus increase $\mathcal{L}_{w}(q_{w|x,y},p;x,y)$.
Importantly, for a truly condition-agnostic $q_{z|x}$, the conditional expectations $\mu_{k}$ will collapse to a shared mean $\mu\coloneqq\mathbb{E}_{p_{x}}\left[\mathbb{E}_{q_{z|x}}\left[z\right]\right]$. In this case $p_{w|y}$ becomes a shared prior across classes, centered at a meaningful point in the latent space, rather than an uninformative one.
As the following proposition establishes, this anchoring of the prior $p_{w|y}$ in the variational distribution $q_{z|x}$ preserves the convex–concave structure of the objective, ensuring that the resulting max-min problem has a unique optimal solution.
Proposition 2.2.
Let $\mathcal{Q}_{z}$ and $\mathcal{Q}_{w}$ be convex parametric families of variational distributions over $z$ and $w$, respectively, and let $\mathcal{G}$ denote a convex set of classifiers such that $g(x)\in[0,1]$ for all $g\in\mathcal{G}$. Assume the latent priors are given by $z\sim p(z)$ and $p(w|y)=\mathcal{N}(\mu_{y},I)$, where $p(z)$ is a continuous strictly positive distribution, and $\mu_{y}=\mathbb{E}_{p_{x|y}}\left[\mathbb{E}_{q_{z|x}}[z]\right]$. Then, under standard regularity conditions (see Appendix B.2.1), there exists a unique saddle point:
$\left(q_{z|x}^{*},q_{w|x,y}^{*},g^{*}_{y\mid z}\right)=\max_{q_{z|x}\in\mathcal{Q}_{z}}\max_{q_{w|x,y}\in\mathcal{Q}_{w}}\min_{g_{y|z}\in\mathcal{G}}\mathcal{L}(q_{z|x},q_{w|x,y},g_{y|z}).$
The proof is provided in Appendix B.2.2.
Specification of variational families
We set both variational families $\mathcal{Q}_{z}$ and $\mathcal{Q}_{w}$ as $d$-dimensional Gaussian distributions with diagonal covariance matrices. Accordingly, each variational distribution is parameterized by a mean vector $\mu\in\mathbb{R}^{d}$ and a vector of variances $\sigma^{2}\in\mathbb{R}^{d}_{+}$ corresponding to the diagonal of the covariance matrix, yielding $\theta=(\mu,\sigma^{2})$.
3 Encoder-decoder model
In order to optimize the objective in Equation 8 with respect to $q_{z|x},q_{w|x,y}$ and $g_{y|z}$ over the dataset $\mathcal{D}$, we introduce an encoder-decoder framework (for illustration see Figure 1(b)) in which two separate reconstructions of $x$ are generated: one, denoted $\hat{x}\sim p_{x|z}$, where $z$ is sampled from the condition-invariant posterior $q_{z|x}$, and the other, denoted $\tilde{x}\sim p_{x|z,w}$, where in addition, $w$ is sampled from the condition-aware posterior $q_{w|x,y}$. The corresponding algorithm is summarized in Algorithm 1.
Condition agnostic representation   An input $x\in\mathcal{X}$ is mapped to the variational parameters $\theta_{z}=(\mu_{z},\sigma^{2}_{z})$ by an encoder neural network $f^{z}_{\phi}:\mathcal{X}\rightarrow\mathbb{R}^{d}\times\mathbb{R}^{d}_{+}$ parametrized by weights $\phi$. A latent encoding $z\sim q_{\theta_{z}}$ is then sampled and mapped to a reconstruction $\hat{x}$ via a decoder neural network $h^{z}_{\psi}:\mathbb{R}^{d}\rightarrow\mathcal{X}$ parametrized by weights $\psi$.
Adversarial classifier   To avoid constructing a complex classifier directly from $z$ to $y$, we leverage the reconstruction $\hat{x}$ from $z$. As a proxy for $g(y|z)$, we instead predict $y$ from $\hat{x}$, allowing the use of a simple logistic regression on $\hat{x}$. Therefore, we define the adversarial classifier as a multinomial logistic regression $g_{\beta}:\hat{\mathcal{X}}\rightarrow[0,1]^{K}$, parametrized by class-specific weight vectors $\beta=\{\beta_{k}\}_{k=1}^{K}$.
While, from an information-theoretic perspective, this substitution weakens the estimation of the cross-entropy term $-\mathbb{E}_{q_{z|x}}\log g(y\mid z)$, since by the data processing inequality, $\hat{x}$ can contain at most the information present in $z$, we observed this to be often preferable from an optimization perspective. Predicting $y$ from $\hat{x}$ helps reduce the variance introduced by sampling $z\sim q_{\theta_{z}}$, and thus acts as a form of regularization that prevents $q_{\theta_{z}}$ from overfitting to noisy penalties produced by the classifier.
Condition aware representation   A labeled input pair $(x,y)\in\mathcal{X}\times\{1,\dots,K\}$ is mapped to the parameters $\theta_{w}=(\mu_{w},\sigma^{2}_{w})$ using an encoder neural network neural-network $f^{w}_{\rho}:\mathcal{X}\times\{1,\dots,K\}\rightarrow\mathbb{R}^{d}\times\mathbb{R}^{d}_{+}$ parametrized by weights $\rho$. A sample $w\sim q_{\theta_{w}}$ is then drawn, and the pair $(z,w)$ is mapped to a reconstruction $\tilde{x}$ via a decoder neural-network $h^{z,w}_{\eta}:\mathbb{R}^{d}\rightarrow\mathcal{X}$, parametrized by weights $\eta$. To compute the prior $p_{w|y}$, we estimate the class-specific mean as $\hat{\mu}_{k}\coloneqq\frac{1}{n_{k}}\sum_{i;y_{i}=k}z_{i}$ where each $z_{i}\sim p(z\mid x_{i})$ is sampled from the encoder given an input $x_{i}$ with label $y_{i}=k$, and $n_{k}$ is the number of training points with the label $y=k$.
Algorithm 1
1:  Input: Data $\mathcal{D}=\{x_{1:n},y_{1:n}\}$, number of training iterations $J$, initial parameters $\phi^{(0)},\psi^{(0)},\rho^{(0)},\eta^{(0)},\beta^{(0)}$, learning rates $\gamma_{1},\gamma_{2}$, weighting terms $\alpha=(\alpha_{1},\alpha_{2})$
2:  for $1\leq j\leq J$ do
3:     Compute $\theta_{z}=f^{z}_{\phi^{z}}(x)$ and $\theta_{w}=f^{w}_{\rho^{(j-1)}}(x,y)$
4:     Sample condition invariant and aware latent variables $z\sim q_{\theta_{z}}$ and $w\sim w_{\theta_{w}}$
5:     Compute reconstructions $\hat{x}=h^{z}_{\psi^{(j-1)}}(z)$ and $\tilde{x}=h^{z,w}_{\eta^{(j-1)}}$
6:     Compute condition prediction $\hat{y}=g_{\beta^{(j-1)}}(\hat{x})$
7:     Update classifier parameters:
$\beta^{(j)}\leftarrow\beta^{(j-1)}-\gamma_{1}\nabla_{\beta}\,\mathcal{L}_{\alpha}(q_{z|x},q_{w|x,y},g_{y|z})\Big{|}_{\phi=\phi^{(j-1)},\psi=\psi^{(j-1)},\rho=\rho^{(j-1)},\eta=\eta^{(j-1)}}$
8:     Update encoder and decoder parameters, $\Omega^{(j)}=\left(\phi^{(j)},\psi^{(j)},\rho^{(j)},\eta^{(j)}\right),$
$\Omega^{(j)}\leftarrow\Omega^{(j-1)}-\gamma_{2}\nabla_{\phi,\psi,\rho,\eta}\,\mathcal{L}_{\alpha}(q_{z|x},q_{w|x,y},g_{y|z})\Big{|}_{\beta=\beta^{(j)}}$
9:  end forReturn: $\phi^{(J)},\psi^{(J)},\rho^{(J)},\eta^{(J)},\beta^{(J)}$
In practice, following standard approaches in VAE-based models, we make two approximations: (i) we use a single-sample Monte Carlo estimate to approximate the expectations in Equation 8, and (ii) instead of directly sampling from $q_{\theta}$, we employ the reparameterization trick to enable differentiable sampling. Specifically, we sample $\epsilon\sim\mathcal{N}(0,I)$ and obtain a sample from $q_{\theta}$ by applying a deterministic transformation of $\epsilon$ based on the variational parameters $\theta$.
4 Comparison to previous approaches
Here we review VAE-based methods for disentangled representation learning, which form the primary basis for comparison with our approach. Broader related literature is discussed in Appendix A.
VAEs (Kingma and Welling, 2014) are generative models that learn latent representations by maximizing the evidence lower bound (ELBO) on the data log-likelihood:
$\mathbb{E}_{q_{z|x}}[\log p(x\mid z)]-\operatorname{D_{\mathrm{KL}}}\left(q_{z\mid x}\ \|\ p_{z}\right)\leq\log p(x),$
where $(x,z)\sim p$, and $z|x\sim q$ is a latent variable inferred from data. VAEs consist of an encoder $q_{z\mid x}$ that maps inputs to latent distributions, and a decoder $p_{x\mid z}$ that reconstructs inputs from latent representations. The learning process frames posterior inference as KL-regularized optimization over a variational family $\mathcal{Q}$ aiming to approximate the posterior $p_{z|x}$ under a typically simple prior $p(z)$. Several VAE extensions were proposed to encourage disentanglement. These are discussed below.
Conditional VAEs   (Sohn et al., 2015) incorporate supervision to the standard VAE model by conditioning both the encoder and decoder on an observed label $y$, yielding the following objective:
$\mathbb{E}_{q_{z|x,y}}[\log p(x\mid z,y)]-\operatorname{D_{\mathrm{KL}}}\left(q(z\mid x,y)\|p(z)\right)\leq\log p(x|y).$
While this allows controlled generation and partial disentanglement between $z$ and $y$, the latent variable $z$ is still inferred from both $x$ and $y$ and thus can encode condition-specific information.
Conditional Subspace VAEs  (CSVAEs) (Klys et al., 2018), explicitly factorize the latent space into a shared component $z$ a label-specific component $w$ (see Supplementary Figure 1(a)). Similarly, their hierarchical extension Beker et al. (2024) introduces an intermediate latent variable between $x$ and $(z,w)$. As in our approach, to encourage disentanglement, CSVAEs introduce an adversarial regularization term that penalizes mutual information between $z$ and $y$, thereby discouraging predictability of $y$ from $z$. They are learned by optimizing the following lower bound on $\log p(x\mid y)$:
[…]
Domain Invariant VAEs (DIVA) and Characteristic-capturing VAEs (CCVAE)   DIVA (Ilse et al., 2020) and CCVAE (Joy et al., 2020), shown in Supplementary Figure 1(b), introduce two latent variables, $z$ and $w$, where $w$ captures label-related features by jointly optimizing a classifier $p(w\mid y)$ jointly alongside the remaining objective. For fully supervised cases, the DIVA model optimizes
$\textstyle\mathbb{E}_{q_{z,w|x}}[\log p(x\mid z,w)]+\mathbb{E}_{q_{w|x}}[\log q(y\mid w)]-\operatorname{D_{\mathrm{KL}}}\left(q_{z|x}\|p_{z}\right)-\operatorname{D_{\mathrm{KL}}}\left(q_{w\mid x}\|p_{w\mid y}\right)\leq\log p(x,y).$
This objective corresponds to the assumptions $x\perp y\mid z,w$ and $z\perp w$. Similarly, CCVAEs optimize the following lower bound on $\log p(x\mid y)$:
$\textstyle\mathbb{E}_{q_{z,w|x,y}}\left[\frac{q(y\mid w)}{q(y\mid x)}\log\frac{p(x\mid z,w)}{q(y\mid w)}\right]-\operatorname{D_{\mathrm{KL}}}\left(q_{z|x}\|p_{z|y}\right)-\operatorname{D_{\mathrm{KL}}}\left(q_{w\mid x}\|p_{w\mid y}\right)+\log q(y\mid x).$
Comparison:   In contrast to our approach, which relies on minimal independence Assumptions (1 -2), CSVAEs, DIVA, and CCVAE incorporate stronger structural assumptions directly into their objectives. Additionally, DIVA and CCVAE lack explicit mechanisms to remove residual dependence between $z$ and $y$, often leading to information leakage into the shared representation.
5 Experiments
We evaluate DISCoVeR with existing methods on both synthetic datasets and real datasets. When applicable, we assess reconstruction quality using negative log-likelihood (NLL), root mean squared error (RMSE), and the absolute deviation from the optimal Bayes classifier on the completed reconstruction, denoted as $\Delta$ – Bayes. We measure disentanglement using a neural estimation Belghazi et al. (2018) of the the mutual information $I(z;w)$. For all experiments, model architectures, hyper-parameters and all other implementation details are provided in Appendix E.
5.1 Simulated data
We begin with controlled synthetic experiments to isolate and visualize disentanglement.
5.1.1 Parametric model
Data generating model: Consider a model where the observed data $x$ is generated as a function of two latent variables $z$ and $w$, and $y$ are binary labels. Assume that the marginal distributions of the latent variables are given by $z\sim\mathcal{N}(0,1)$ and $w\sim\mathcal{N}(0,1)$, and that the data $x$ is generated as the sum of the two latent variables: $x=z+w$. Since $z$ and $w$ are both drawn from $\mathcal{N}(0,1)$, it follows that $x\sim\mathcal{N}(0,2)$. Finally, assume that the binary label is determined by the sign of $w$: $y=1$ if $w>0$, and $y=0$ otherwise.
Optimal disentanglement: Given that $z$ and $w$ are independent and $x=z+w$, we have that $p(z\mid x)=\mathcal{N}\!\left(z;\frac{x}{2},\frac{1}{2}\right)$. Hence, given $x$, the best estimate for $z$ is $\frac{x}{2}$. Note that when ignoring the label $y$, the conditional distribution $p(w\mid x)$ is $p(w\mid x)=\mathcal{N}\!\left(w;\frac{x}{2},\frac{1}{2}\right)$. However, the observation of $y$ (which indicates whether $w$ is positive or negative) truncates this distribution:
$p(w\mid x,y=1)=\frac{\mathcal{N}\left(w;\frac{x}{2},\frac{1}{2}\right)}{\Phi\left(\frac{x}{\sqrt{2}}\right)},\qquad p(w\mid x,y=0)=\frac{\mathcal{N}\left(w;\frac{x}{2},\frac{1}{2}\right)}{1-\Phi\left(\frac{x}{\sqrt{2}}\right)}.$ (11)
Results: Table 1 shows that DISCoVeR (ours) best approximates the analytic posteriors, resulting in the lowest deviation from the optimal Bayes classifier and the best reconstruction.
Table 1: Parametric model results: DISCoVeR (ours) outperforms all competitors across all metrics.
NLL $\downarrow$ $\operatorname{D_{\mathrm{KL}}}(q_{z\mid x}\;||\>p_{z\mid x})\downarrow$ $\operatorname{D_{\mathrm{KL}}}(q_{w\mid x}\;||\>p_{w\mid x})\downarrow$ $\Delta$ – Bayes $\downarrow$
CSVAE - N.A. $1.810\pm 0.016$ $6.65\pm 3.46$ $23.61\pm 0.36$ $24.83\pm 0.04$
CSVAE $1.786\pm 0.022$ $2.85\pm 1.11$ $23.98\pm 4.36$ $24.33\pm 1.28$
HCSVAE - N.A. $1.784\pm 0.010$ $4.01\pm 0.07$ $25.82\pm 0.38$ $24.99\pm 0.01$
HCSVAE $1.770\pm 0.004$ $3.99\pm 0.09$ $26.25\pm 0.59$ $24.99\pm 0.01$
DIVA $1.788\pm 0.008$ $3.21\pm 1.52$ $12.88\pm 3.31$ $3.51\pm 0.32$
CCVAE $1.785\pm 0.006$ $1.77\pm 0.81$ $12.95\pm 3.35$ $3.57\pm 0.15$
DISCoVeR (ours) $\mathbf{1.769\pm 0.003}$ $\mathbf{0.17\pm 0.01}$ $\mathbf{10.10\pm 0.73}$ $\mathbf{0.1\pm 0.28}$
5.1.2 Noisy Swiss roll
Dataset:   We use a noisy variant of the labeled Swiss Roll dataset (Marsland, 2014; Klys et al., 2018), generating $n=20,000$ and assigning binary labels based on a lengthwise split, with labels flipped at rate $\rho$. The common geometry (its projection along the 2D plane) remains intact, he conditional structure along the third axis becomes noisy. Figure 2A illustrates the setup.
Table 2: Noisy Swiss roll results: DISCoVeR  yields lowest deviation from optimal-Bayes, while maintaining low latent leakage and high reconstruction accuracy.
$I(z;w)\downarrow$ NLL $\downarrow$ $\Delta$ – Bayes $\downarrow$
CSVAE - N.A. $0.011$ $3.303\pm 0.003$ $23.88\pm 12.02$
CSVAE $\mathbf{0.002}$ $\mathbf{3.302\pm 0.003}$ $17.99\pm 14.58$
HCSVAE - N.A. $0.062$ $\mathbf{3.302\pm 0.002}$ $30.00\pm 0.00$
HCSVAE $0.003$ $\mathbf{3.302\pm 0.002}$ $30.00\pm 0.00$
DIVA $0.423$ $\mathbf{3.302\pm 0.003}$ $1.88\pm 1.05$
CCVAE $0.254$ $\mathbf{3.302\pm 0.002}$ $2.21\pm 0.84$
DISCoVeR (Ours) $0.004$ $\mathbf{3.302\pm 0.002}$ $\mathbf{1.14\pm 0.21}$
Optimal disentanglement:   Since the Swiss Roll is sliced at the center and label noise is applied uniformly, marginalizing over labels yields a symmetric spiral centered along the roll—i.e., the marginal posterior $p(z\mid x)$ is label-invariant. In contrast, the conditional component retains a noisy but informative signal, with a uniform noise level of $\rho=0.3$. As a result, the Bayes optimal classifier trained on any realistic representation is upper-bounded at 70% accuracy.
Results:   Figure 2 presents qualitative and quantitative results, showing that DISCoVeR both models the label noise accurately and effectively disentangles shared and condition-specific structure. Notably, DISCoVeR captures the marginal data distribution, successfully recovering the expected spiral pattern, as shown in Figure 2B (right). Additionally, the results in Table 2 show that DISCoVeR achieves the lowest deviation from the optimal Bayes classifier and minimal information leakage between latent variables, while preserving reconstruction quality.
Figure 2: A: Noisy labeled Swiss Roll dataset. B (left): Reconstruction and conditional embedding from DISCoVeR. B (right): Marginal reconstruction from the shared embedding, recovering the spiral structure. C: Comparison of reconstruction and conditional embeddings across models.
5.2 Real data
5.2.1 Noisy colored MNIST
Dataset:   We use a modified MNIST (Deng, 2012) dataset constructed from $60,000$ duplicated images: in one copy we remove the red channel ($y=0$) and in the other we remove the green channel ($y=1$). The digit shape remains intact following the blue channel. Label noise is introduced by flipping labels with probability $\rho\in\{0,0.1,0.2,0.3,0.4\}$.
Optimal disentanglement:   Since the colored images are generated in equal proportions, the marginal reconstruction for retains a single color (see Figure 3).
Results:   We evaluate marginal coloring reconstruction by DISCoVeR and previous methods. Under no label noise ($\rho=0$), all methods perform similarly (Supplementary Figure 3). At all non-zero noise levels, DISCoVeR consistently outperforms competitors, whose marginal reconstructions are averaged over class-conditioned outputs. Metrics for $\rho=0.3$ are shown in Supplementary Table 2, with results for other noise levels in Supplementary Figure 4.
Figure 3: Colored MNIST results for noise level $\rho=0.3$. DISCoVeR is the only model that consistently reconstructs the correct constant color matching the true marginal $p(x|B)$.
5.2.2 CelebA-Glasses
Dataset:   We use all CelebA (Liu et al., 2015) images labeled with eyeglasses attribute ($y=1$), and twice as many randomly sampled images without ($y=0$), totaling $n=35,712$ samples.
Results:   Figure 4 shows that DISCoVeR accurately reconstructs input images while producing shared embeddings that marginalize over the eyeglasses attribute, consistently adding "pseudo-glasses" to all samples. Competing methods are shown in Supplementary Figure 5, with quantitative results in Supplementary Table 3. While reconstruction quality is comparable across methods, DISCoVeR achieves significantly better disentanglement.
Figure 4: CelebA-Eyeglasses results. Top: Original images with and without eyeglasses. Middle: Full reconstructions by DISCoVeR. Bottom: Reconstructions solely from common embeddings $z$.
5.2.3 Single cell RNA-sequencing (scRNA-seq) data from Lupus dataset
Dataset:   We analyze single-cell RNA sequencing from $n=13,999$ peripheral blood mononuclear cells (PBMCs) collected from 8 lupus patients under two conditions: 7,451 cells control ($y=0$), and 6,548 IFN-$\beta$ stimulation cells. IFN-$\beta$ stimulation induces notable shifts in gene expression, visible in the UMAP embedding in Figure 5B (left).
Figure 5: A, B (left): UMAP embeddings of gene counts from the IFN-$\beta$ stimulation dataset. A, B (middle): Common embedding $z$ preserves cell type-specific differences, while correcting for the stimulation effect. A, B (right): Conditional embedding $w$ captures the stimulation effect. C: Comparison of original and reconstructed log-mean (top) and log-variance (bottom) across cells for the two different reconstructions offered by DISCoVeR.
Results:   Supplementary Table 5 shows that DISCoVeR effectively achieves the desired behavior with strong empirical performance, where only cell type information is captured in $z$ (Figure 5A, middle) while the effects of IFN-$\beta$ stimulation are wholly represented in $w$ (Figure 5B, right). Other approaches either (1) achieve mixing in the $z$ space, but compromise on keeping cell types separated or (2) leak information about stimulation into the $z$ space (Supplementary Figure 6).
Facilitating interpretability:   By enabling marginalized reconstructions, DISCoVeR provides a direct link between shared embeddings and gene expression, offering clearer insight into the effects of IFN-$\beta$ stimulation, unlike other methods. In Figure 5C, comparing variance across marginal and full reconstructions accurately recovers gene-level differences associated with IFN-$\beta$ stimulation, including ISG15, FTL, CCL8, CXCL10, CXCL11, APOBEC3A, IL1RN, IFITM3 and RSAD2.
6 Conclusion
In this work we introduced a variational framework for disentangled representation learning that explicitly separates condition-invariant and condition-aware factors, while avoiding overly restrictive independence assumptions. Unlike prior work, DISCoVeR incorporates two reconstruction paths—one based solely on the shared latent variable $z$, and the other on the full pair $(z,w)$. Our model simultaneously learns informative shared representations, and captures structured variation across conditions. Experimental results demonstrate that DISCoVeR achieves strong reconstruction, minimal information leakage, and accurate modeling of conditional effects—consistently outperforming existing methods.
Our model assumes that the condition labels are fully observed and discrete. We leave extensions of our framework to semi-supervised or continuous condition settings for future work.
Appendix A Additional related work
A.1 Domain Generalization
The task of representation disentanglement is closely related to the field of domain generalization [Muandet et al., 2013], which assumes limited or no access to target domain samples and aims to learn representations that can be readily adapted, often via transfer learning, to new, unseen domains.
As noted by Wang et al. [2019], existing methods in domain generalization can be broadly categorized into two main approaches: (i) approaches for reducing the inter-domain differences, often by using adversarial techniques [Ghifary et al., 2015, Wang et al., 2017, Motiian et al., 2017, Li et al., 2018, Carlucci et al., 2019, Li et al., 2018, Wang et al., 2019, Akuzawa et al., 2020, Zhu et al., 2022, Gokhale et al., 2023, Dayal et al., 2023, Cheng et al., 2023, Chen et al., 2024], and (ii) Approaches that construct an ensemble of domain-specific models, and then fuse their representations to form a unified, domain-agnostic representation [Ding and Fu, 2017, Mancini et al., 2018, Zhou et al., 2021, Muhammad et al., 2024].
Additional strategies for domain generalization include contrastive learning approaches [Kim et al., 2021], methods based on distribution alignment via metrics [Muandet et al., 2013, Sun and Saenko, 2016], and techniques utilizing custom network architectures, for instance by incorporating domain-specific adapters between shared layers [Rebuffi et al., 2017, 2018, Li and Vasconcelos, 2019, Omi et al., 2022].
The primary distinction between these methods and ours lies in the explicit probabilistic modeling and disentanglement of domain-invariant and domain-specific factors. Whereas prior approaches typically focus on aligning domains through adversarial training or fusing multiple domain-specific predictors, our method constructs a structured latent space, decomposed into a shared representation $z$, capturing domain-invariant information, and a conditional component $w$, which encodes domain-specific variability. This factorization is learned through a tailored variational objective involving an adversarial penalty and two reconstructions —one based on $z$ alone, and another on the full latent pair $(z,w)$, thereby promoting both interpretability and a clean separation of shared and domain-aware features.
A.2 Out of distribution generalization
A.2.1 Enviroment balancing methods
The field of out-of-distribution (OOD) generalization emerged from foundational work on causality and invariance across training environments [Peters et al., 2016, 2017]. The central assumption is that each environment exhibits distinct spurious correlations between features and labels; therefore, robust generalization requires models to focus on invariant relationships that hold across environments. To address this distribution shift, many recent approaches adopt a regularized empirical risk minimization framework:
$\min_{\theta}\quad\smashoperator[]{\sum_{e\in E_{\text{train}}}^{}}\ell^{e}(f_{\theta})+\lambda R(f_{\theta},E_{\text{train}}),$ (12)
where the regularizer $R$ encourages representations that are stable across environments. Among these, Invariant Risk Minimization (IRM) [Arjovsky et al., 2019] enforces that a single classifier remains optimal across all environments, Variance Risk Extrapolation (VarREx) [Krueger et al., 2021] promotes robustness by minimizing the variance of losses across environments, and CLOvE [Wald et al., 2021] takes a calibration-theoretic perspective, penalizing discrepancies between predicted confidence and correctness across environments.
While these methods focus on enforcing predictive invariance across environments through regularization, our approach instead explicitly enforces conditional independence between the shared latent variable $z$ and an environment-aware variable $w$.
A.2.2 Distributionally robust methods
An alternative line of work for handling distribution shifts is Distributionally Robust Optimization (DRO) [Ben-Tal et al., 2013, Duchi et al., 2021, Duchi and Namkoong, 2021, Wei et al., 2023], which avoids assuming a fixed data-generating distribution. Instead, DRO methods optimize performance under the worst-case scenario over a family of plausible distributions. A prominent variant, known as group DRO [Sagawa et al., 2019, Piratla et al., 2021], introduces group-level structure that may correlate with spurious features, potentially leading to biased predictions. In settings where group labels are not directly observed, several strategies have been proposed, including reweighting high-loss examples [Liu et al., 2021] and balancing class-group combinations through data sub-sampling [Idrissi et al., 2022].
However, these approaches assume that the label space remains fixed between training and test time, limiting their applicability in adaptation to new domains, environments or conditions.
A.3 Zero-shot learning
Zero-shot learning systems [Fei-Fei et al., 2006, Larochelle et al., 2008] aim to classify instances from novel, previously unseen classes at test time. In contrast to the out-of-distribution (OOD) generalization setting, these approaches typically do not assume the presence or structure of a distribution shift. Instead, a common strategy is to learn data representations that capture class-agnostic similarity, enabling the model to determine whether two instances belong to the same class without requiring knowledge of the class identity itself. Such methods include contrastive-learning [Hadsell et al., 2006], siamese neural networks [Koch et al., 2015], triplet networks [Hoffer and Ailon, 2015], and other more recent variations [Oh Song et al., 2016, Sohn, 2016, Wu et al., 2017, Yuan et al., 2019]. Recent work has begun to address the impact of class distribution shifts in zero-shot settings. For instance, Slavutsky and Benjamini [2024] integrate environment-based regularization—motivated by OOD generalization—with zero-shot learning by simulating distribution shifts through hierarchical sampling, enabling the model to learn representations that are robust to shifts in class distributions.
While this line of work shares our motivation of improving robustness under unseen conditions, it primarily addresses the problem of class-level generalization through similarity-based learning, rather than explicitly modeling and disentangling the latent factors—such as domain or environment—that drive distributional variation across tasks.
Appendix B Proofs
B.1 Proof of Proposition 2.1
Proof.
$\operatorname{ELBO}(q,p;x,y)-\mathcal{L}_{w}(q_{w|x,y},p;x,y)$ (13)
$=\left[\log p(x\mid y)-\operatorname{D_{\mathrm{KL}}}(q_{w|x,y}\,\|\,p_{w|x,y})\right]-\left[\log p(x\mid y)-\operatorname{D_{\mathrm{KL}}}\left(q_{z|x}q_{w|x,y}\,\|\,p_{z,w|x,y}\right)\right]$ (14)
$=\operatorname{D_{\mathrm{KL}}}\left(q_{z|x}q_{w|x,y}\,\|\,p_{z,w|x,y}\right)-\operatorname{D_{\mathrm{KL}}}\left(q_{w|x,y}\,\|\,p_{w|x,y}\right)$ (15)
$=\mathbb{E}_{q_{w|x,y}}\left[\mathbb{E}_{q_{z|x}}\left[\log q(z\mid x)+\log q(w|x,y)-\log p(z,w\mid x,y)\right]\right]$ (16)
$\phantom{=}-\mathbb{E}_{q_{w|x,y}}\left[\log q(w|x,y)-\log p(w|x,y)\right]$ (17)
$=\mathbb{E}_{q_{w|x,y}}\left[\mathbb{E}_{q_{z|x}}\left[\log q(z\mid x)-\log p(z,w\mid x,y)+\log p(w|x,y)\right]\right]$ (18)
$=\mathbb{E}_{q_{z|x}}\left[\log q(z\mid x)-\log p(z\mid w,x,y)\right]$ (19)
$=\operatorname{D_{\mathrm{KL}}}\left(q_{z|x}\,\|\,p_{z|w,x,y}\right).$ (20)
B.2 Game equilibrium
B.2.1 Regularity conditions
To ensure that expectations and KL-terms in the game objective $\mathcal{L}(q_{z|x},q_{w|x,y},g_{y|z})$ render the functionals strictly concave in $q_{z|x}$ , strictly concave in $q_{w|x,y}$, and strictly convex in $g$, the following regularity conditions are required:
  1. 1.
    The likelihoods $p(x|z),p(x|z,w),p(y|x)$ are strictly positive, continuous densities.
  2. 2.
    The variational families $Q_{z}$ and $Q_{w}$, and the set of achievable classifiers $\mathcal{G}$ are non-empty, convex and compact.
  3. 3.
    $\log p(x|z,w)$ and $\log g(y|x)$ are integrable.
B.2.2 Proof of Proposition 2.2
Proof.
Since $\mathcal{L}_{z}(q_{z|x},p;x)$ is the standard ELBO objective, we have that
$\mathcal{L}_{z}(q_{z|x},p;x)=\log p(x)-\operatorname{D_{\mathrm{KL}}}\left(q_{z|x}\,\|\,p_{z|x}\right).$ (21)
Similarly, we have that
$\mathcal{L}_{w}(q_{w|x,y},p;x,y)$ $=\log p(x\mid y)-\operatorname{D_{\mathrm{KL}}}\left(q_{z|x}q_{w|x,y}\,\|\,p_{z,w|x,y}\right).$ (22)
Thus,
$\mathcal{L}(q_{z|x},q_{w|x,y},g_{y|z})=\mathbb{E}_{p_{x,y}}$ $\left[\log p(x)-\operatorname{D_{\mathrm{KL}}}\left(q_{z|x}\,\|\,p_{z|x}\right)\right.$ (23)
$+\log p(x\mid y)-\operatorname{D_{\mathrm{KL}}}\left(q_{z|x}q_{w|x,y}\,\|\,p_{z,w|x,y}\right)$ (24)
$\left.-\mathbb{E}_{q_{z|x}}\log g(y\mid z)\right].$ (25)
For fixed $q_{z|x}$, the adversarial classifier minimizes:
$-\mathbb{E}_{p_{x,y}}\mathbb{E}_{q_{z|x}}\log g(y\mid z),$ (26)
which is the population cross-entropy and is strictly convex in $g(y|z)$, and thus has a unique solution.
It remains to show that the terms in the objective function that depend on $q_{z|x}$ and $q_{w|x,y}$, are strictly concave in each argument when the others are held fixed.
Focusing on the terms dependent on $q_{w|x,y}$ first, define
$\ell_{w}\coloneqq$ $-\operatorname{D_{\mathrm{KL}}}\left(q_{z|x}q_{w|x,y}\,\|\,p_{z,w|x,y}\right)$ (27)
[…]
$=$ $H(q_{z\mid x})+H(q_{w\mid x,y})+\mathbb{E}_{q_{z|x}}\mathbb{E}_{q_{w|x,y}}\log p(z,w\mid x,y).$ (30)
Note that
$\mathbb{E}_{q_{z|x}}\mathbb{E}_{q_{w|x,y}}\log p(z,w\mid x,y)$ (31)
is linear in $q_{w|x,y}$, and since $H(q_{w\mid x,y})$ is strictly concave in $q_{w|x,y}$, we have that $\mathbb{E}_{p_{x,y}}[\ell_{w}]$ is strictly concave in $q_{w|x,y}$.
Similarly, define
$\ell_{z}$ $\coloneqq-\operatorname{D_{\mathrm{KL}}}\left(q_{z|x}\,\|\,p_{z|x}\right)-\operatorname{D_{\mathrm{KL}}}\left(q_{z|x}q_{w|x,y}\,\|\,p_{z,w|x,y}\right).$ (32)
By convexity of KL divergence in its first argument, $-\operatorname{D_{\mathrm{KL}}}\left(q_{z|x}\,\|\,p_{z|x}\right)$ is strictly concave in $q_{z\mid x}$.
Focusing on the second KL term, from Equation 30 we have that
$-\operatorname{D_{\mathrm{KL}}}\left(q_{z|x}q_{w|x,y}\,\|\,p_{z,w|x,y}\right)=H(q_{z\mid x})+H(q_{w\mid x,y})+\mathbb{E}_{q_{z|x}}\mathbb{E}_{q_{w|x,y}}\log p(z,w\mid x,y),$ (33)
where $H(q_{z\mid x})$ is strictly concave in $q_{z|x}$.
Recall that we assumed that $p(w|y)$ depends on $q(z|x)$. Under our model
$p(x,y,z,w)=p(y)p(w\mid y)p(z)p(x\mid z,w),$ (34)
yielding
$p\left(z,w,\mid x,y\right)=p(w\mid y)\frac{p(z)p(x\mid z,w)}{p(x\mid y)}.$ (35)
Hence,
$\mathbb{E}_{q_{z|x}}\mathbb{E}_{q_{w|x,y}}\log p(z,w\mid x,y)=\mathbb{E}_{q_{z|x}}\mathbb{E}_{q_{w|x,y}}\left[\log p(w\mid y)+\log\frac{p(z)p(x\mid z,w)}{p(x\mid y)}\right],$
where $p(w\mid y)=\mathcal{N}\left(w;\mu_{y},I\right)$ with $\mu_{y}=\mathbb{E}_{p_{x|y}}\left[\mathbb{E}_{q_{z|x}}[z]\right]$. Therefore,
$\mathbb{E}_{q_{z|x}}\mathbb{E}_{q_{w|x,y}}\left[\log p(w\mid y)\right]=-\frac{1}{2}\left[d\log(2\pi)+\mathbb{E}_{q_{z|x}}\mathbb{E}_{q_{w|x,y}}\left\|w-\mu_{y}\right\|^{2}\right]$ (36)
where $-\left\|w-\mu_{y}\right\|^{2}$ is a quadratic form in $\mu_{y}$, which is linear in $q_{z|x}$, and thus $\mathbb{E}_{q_{z|x}}\mathbb{E}_{q_{w|x,y}}\left[\log p(w\mid y)\right]$ is strictly concave in $q_{z|x}$. Hence, $-\operatorname{D_{\mathrm{KL}}}\left(q_{z|x}q_{w|x,y}\,\|\,p_{z,w|x,y}\right)$ is strictly concave in $q_{z|x}$, and thus so is $\mathbb{E}_{p_{x,y}}[\ell_{z}]$. ∎
Appendix C Supplementary Figures
$y$$\,w\mid x,y\,$$\,z\mid x\,$$\hat{y}$$x$$\,\tilde{x}\mid z,w\,$
$\,w\mid x\,$$\,z\mid x\,$$\hat{y}$$x$$\,\tilde{x}\mid z,w\,$
Supplementary Figure 1: Encoder-decoder structures for previous approaches. (a) CSVAE. (b) DIVA - CCVAE.
Supplementary Figure 2: Comparison of approximate variational posteriors against the true posterior for latent variables $z,w$ for different values of $x$ with $y=0$ (top) and $y=1$ (bottom).
Supplementary Figure 3: Colored MNIST results for no noise.
Supplementary Figure 4: Colored MNIST visual results across the remaining noise levels.
Supplementary Figure 5: Reconstruction performance for other models on the CelebA-Glasses dataset. Top: Original samples from the data. Bottom: Reconstructions by the given model.
Supplementary Figure 6: Embeddings obtained by other models on the Kang dataset. For each block, top (resp. bottom) rows are $z$ (resp. $w$) embeddings, while left (resp. right) columns are colored by cell type (resp. stimulation).
Appendix D Supplementary Tables
Supplementary Table 1: RMSE for the Colored MNIST dataset without any label noise.
Marginal RMSE ($\rho=0$) $\downarrow$
CSVAE - No Adv. $\mathbf{0.064\pm 0.002}$
CSVAE $0.079\pm 0.008$
HCSVAE - No Adv. $0.094\pm 0.004$
HCSVAE $0.079\pm 0.030$
DIVA $0.065\pm 0.005$
CCVAE $0.065\pm 0.006$
DISCoVeR (Ours) $\mathbf{0.064\pm 0.000}$
Supplementary Table 2: RMSE calculated between the estimated and true marginal across different levels of label noise on the Colored MNIST dataset. $p$ defines label flip probability. Bold denotes best performance.
Marginal RMSE $\downarrow$
$p=0.1$ $p=0.2$ $p=0.3$ $p=0.4$
CSVAE - No Adv. $0.141\pm 0.002$ $0.141\pm 0.003$ $0.142\pm 0.002$ $0.143\pm 0.002$
CSVAE $0.135\pm 0.022$ $0.152\pm 0.018$ $0.181\pm 0.007$ $0.173\pm 0.008$
HCSVAE - No Adv. $0.150\pm 0.001$ $0.150\pm 0.000$ $0.151\pm 0.000$ $0.151\pm 0.001$
HCSVAE $0.139\pm 0.003$ $0.141\pm 0.001$ $0.141\pm 0.001$ $0.141\pm 0.001$
DIVA $0.115\pm 0.011$ $0.102\pm 0.013$ $0.106\pm 0.010$ $0.113\pm 0.014$
CCVAE $0.092\pm 0.002$ $0.103\pm 0.014$ $0.099\pm 0.011$ $0.092\pm 0.005$
DISCoVeR (Ours) $\mathbf{0.073\pm 0.001}$ $\mathbf{0.083\pm 0.004}$ $\mathbf{0.087\pm 0.002}$ $\mathbf{0.087\pm 0.001}$
Supplementary Table 3: Model performances on the CelebA-Glasses dataset. Bold denotes best performance.
$I(z;w)\downarrow$ NLL ($\downarrow$)
CSVAE - No Adv. $0.048\pm 0.014$ $137.522\pm 0.155$
CSVAE $0.079\pm 0.029$ $145.989\pm 0.336$
HCSVAE - No Adv. $0.055\pm 0.012$ $131.813\pm 0.21$
HCSVAE $0.055\pm 0.014$ $137.319\pm 0.265$
DIVA $0.188\pm 0.028$ $143.528\pm 0.02$
CCVAE $0.083\pm 0.022$ $\mathbf{131.764\pm 0.006}$
DISCoVeR (Ours) $\mathbf{0.030\pm 0.011}$ $135.677\pm 0.007$
DISCoVeR - Common (Ours) $374.114\pm 0.05$
Appendix E Impelementation Details
E.1 Considerations and reproducibility
We run all experiments on a single H100 GPU. Reported means and standard deviations for tables are conducted over 10 repetitions of the experiment with different random seeds. All models are trained using the AdamW [Loshchilov and Hutter, 2019] optimizer until validation loss stops decreasing for 50 epochs. Wherever provided, we use mutual information neural estimation (MINE, Belghazi et al. [2018]) to obtain mutual information estimates. For Naive Bayes classifiers, we use the implementation provided by scikit-learn [Pedregosa et al., 2011]. To use ideal hyperparameters for each method, we consult the original implementation whenever possible, and conduct a simple grid-search to produce originally described model behavior. Implementations of all methods compared in this study, including DISCoVeR, as well as code to reproduce our results, can be found at https://github.com/Computational-Morphogenomics-Group/DISCoVeR. Models compared in the study admit a weighting term for each term in the loss function, of which most are shared across different approaches. We use the following shorthands for each of the terms:
[…]
$\operatorname{D_{\mathrm{KL}}}(Z)\rightarrow\operatorname{D_{\mathrm{KL}}}(q_{z|x}\,\|\,p_{z})$
$\operatorname{D_{\mathrm{KL}}}(W)\rightarrow\operatorname{D_{\mathrm{KL}}}\left(q_{w|x,y}\,\|\,p_{w\mid y}\right)$
[…]
Below, we provide additional details for the hyperparameters used in each experiment, and any other external resources used to obtain the corresponding sections’ results.
E.1.1 Parametric Model
For the parametric model, we consider $z,w\in\mathbb{R}$ and use multi-layer perceptrons (MLPs) with $n_{hidden}=2,d_{hidden}=8$ to parameterize approximate posteriors, the generative model and classifiers. For all models, we use learning rate $\gamma=0.001$. A more detailed table of model-specific loss weights is provided in Supplementary Table 4.
Supplementary Table 4: Loss weights for the parametric model experiment.
Rec. $\operatorname{D_{\mathrm{KL}}}(Z)$ $\operatorname{D_{\mathrm{KL}}}(W)$ Adv. Class. Rec. - ($Z$)
CSVAE - No Adv. 1 1 1
CSVAE 2.5 1 0.5 20
HCSVAE - No Adv. 1 1 0.5
HCSVAE 2.5 1 0.5 20
DIVA 1 1 1 1
CCVAE 1 1 1 1
DISCoVeR (Ours) 0.7 0.7 0.2 0.8 0.3
Supplementary Table 5: K-Means NMI for embeddings across stimulation ($y$) and cell type (common structure).
$w$ Stimulation ($\uparrow$) $z$ Cell Type ($\uparrow$) $z$ Stimulation ($\downarrow$)
CSVAE - No Adv. $0.949\pm 0.003$ $0.702\pm 0.015$ $0.187\pm 0.0$
CSVAE $0.939\pm 0.002$ $0.406\pm 0.001$ $\mathbf{0.002\pm 0.0}$
HCSVAE - No Adv. $0.933\pm 0.006$ $0.628\pm 0.016$ $0.091\pm 0.002$
HCSVAE $0.931\pm 0.005$ $0.433\pm 0.001$ $0.003\pm 0.0$
DIVA $0.801\pm 0.0$ $0.628\pm 0.011$ $0.056\pm 0.0$
CCVAE $0.604\pm 0.0$ $0.683\pm 0.016$ $0.103\pm 0.0$
DISCoVeR (Ours) $\mathbf{0.946\pm 0.002}$ $\mathbf{0.688\pm 0.0}$ $\mathbf{0.002\pm 0.0}$
E.1.2 Noisy Swiss Roll
For this experiment, we consider $z,w\in\mathbb{R}^{2}$ and use MLPs with $n_{hidden}=2,d_{hidden}=128$ to parameterize approximate posteriors, the generative model and classifiers. For all models, we use learning rate $\gamma=0.001$. A more detailed table of model-specific hyperparameters is provided in Supplementary Table 6.
Supplementary Table 6: Loss weights for the noisy Swiss roll experiment.
Rec. $\operatorname{D_{\mathrm{KL}}}(Z)$ $\operatorname{D_{\mathrm{KL}}}(W)$ Adv. Class. Rec. - ($Z$)
CSVAE - No Adv. 20 0.2 1
CSVAE 20 0.2 1 50
HCSVAE - No Adv. 20 0.2 1
HCSVAE 20 0.5 1 50
DIVA 20 0.2 0.2 1
CCVAE 20 0.2 0.2 1
DISCoVeR (Ours) 0.9 0.2 0.2 8 0.1
E.1.3 Noisy Colored MNIST
For this experiment, we consider $z\in\mathbb{R}^{20}$, $w\in\mathbb{R}^{2}$ and use convolutional neural networks (CNNs) to parameterize approximate posteriors and the generative model. For this example, DISCoVeR can support $z,w$ with different sizes, by parameterizing $p(w\mid y)$ through neural networks. For all models, we use learning rate $\gamma=0.0001$. We detail the architectures and model-specific hyperparameters in Supplementary Tables 7 - 10. All other neural networks are formulated as MLPs with $n_{hidden}=2,d_{hidden}=4096$.
Supplementary Table 7: Image encoder architecture for noisy colored MNIST. Parameters for Conv2d are input / output channels. Parameters for MaxPool2D are kernel size and stride. Parameter for the linear layer is the output size. For variances, outputs are passed through an additional Softplus layer to ensure non-negativity.
Block Details
1 Conv2d(3,32) + BatchNorm2D + ReLU
2 Conv2d(32,32) + BatchNorm2D + ReLU + MaxPool2D(2,2)
3 Conv2d(32,64) + BatchNorm2D + ReLU + MaxPool2D(2,2)
4 Conv2d(64,128) + BatchNorm2D + ReLU + MaxPool2D(2,2)
5 Linear(4096) + BatchNorm1D + ReLu
6 Linear(4096) + BatchNorm1D + ReLu
7 Linear($d_{latent}$)
Supplementary Table 8: Image decoder architecture for noisy colored MNIST. Parameters for Conv2d are input / output channels. Parameters for MaxPool2D are kernel size and stride. Parameter for the linear layer is the output size.
Block Details
1 Linear(4096) + BatchNorm1D + ReLu
2 Linear(4096) + BatchNorm1D + ReLu
3 Linear(1152) + Unflatten
4 Upsample(2) + Conv2d(128, 64) + BatchNorm2D + ReLU
5 Upsample(2) + Conv2d(64, 32) + BatchNorm2D + ReLU
6 Upsample(2) + Conv2d(32, 32) + BatchNorm2D + ReLU
7 Conv2d(32, 3) + Sigmoid
Supplementary Table 9: Latent classifier architecture for noisy colored MNIST. Outputs parameterize logits of class probabilities.
Block Details
1 Linear(4096) + BatchNorm1D + ReLu
2 Linear(4096) + BatchNorm1D + ReLu
3 Linear(2)
Supplementary Table 10: Loss weights for the noisy colored MNIST experiment.
Rec. $\operatorname{D_{\mathrm{KL}}}(Z)$ $\operatorname{D_{\mathrm{KL}}}(W)$ Adv. Class. Rec. - ($Z$)
CSVAE - No Adv. 1 0.0001 1
CSVAE 1 0.0001 1 1
HCSVAE - No Adv. 1000 0.0001 1
HCSVAE 10000 0.0001 1 1
DIVA 1 0.0001 0.0001 1
CCVAE 1 0.0001 0.0001 1
DISCoVeR (Ours) 0.5 0.0001 0.0001 0.1 0.5
E.1.4 CelebA-Glasses
Motivated by the previous application of Klys et al. [2018], our choices follow those outlined in Larsen et al. [2016]. We provide a detailed table of model-specific hyperparameters in Supplementary Table 11:
Supplementary Table 11: Loss weights for the CelebA-Glasses experiment.
Rec. $\operatorname{D_{\mathrm{KL}}}(Z)$ $\operatorname{D_{\mathrm{KL}}}(W)$ Adv. Class. Rec. - ($Z$)
CSVAE - No Adv. 1 0.0001 1
CSVAE 1000 0.0001 1 1
HCSVAE - No Adv. 1000 0.0001 1
HCSVAE 10000 0.0001 1 1
DIVA 100000 0.0001 0.0001 1
CCVAE 100000 0.0001 0.0001 1
DISCoVeR (Ours) 1000000 0.0001 0.0001 2000 100000
E.1.5 scRNA-Seq
Following on the previous applications by Lopez et al. [2018], we use $z\in\mathbb{R}^{10}$, $w\in\mathbb{R}^{2}$. For DISCoVeR  we match the dimensions and use $w\in\mathbb{R}^{10}$ to avoid parameterizing the prior $p(w\mid z,y)$ with an additional neural network. We use MLPs with $n_{hidden}=1,d_{hidden}=128$ to parameterize approximate posteriors, the generative model and classifiers. We calculate K-Means NMI through scikit-learn [Pedregosa et al., 2011] by calling the normalized_mutual_info_score function with the original labels and the clusterings obtained by running KMeans on (1) the entire latent embedding and (2) single dimensions of the embedding and report the highest score. A more detailed table of model-specific hyperparameters is provided in Supplementary Table 12:
Supplementary Table 12: Loss weights for the scRNA-seq experiment.
Rec. $\operatorname{D_{\mathrm{KL}}}(Z)$ $\operatorname{D_{\mathrm{KL}}}(W)$ Adv. Class. Rec. - ($Z$)
CSVAE - No Adv. 1 0.0001 1
CSVAE 1 0.0001 1 100
HCSVAE - No Adv. 1 0.0001 1
HCSVAE 1 0.0001 1 100
DIVA 1 0.0001 0.0001 1
CCVAE 1 0.0001 0.0001 1
DISCoVeR (Ours) 0.9 0.0001 0.0001 100 0.1
E.2 Summary of the scVI generative model for 5.2.3
Given batch key $b$ and $G$ genes, the generative model of scVI for a single cell $x_{i}\in\mathbb{N}^{G}$ is formulated as:
$z_{i}\sim\mathcal{N}(0,1)$
$\rho_{i}=f_{\theta}(z_{i},b_{i})$
$\pi_{ig}=h_{\phi}^{g}(z_{i},b_{i})$
$x_{ig}\sim\text{ZINB}(l_{i}\rho_{i},\theta_{g},\pi_{ig})$
Here, $g$ indexes genes, $l_{i}=\sum_{g}x_{ig}$ denotes the total number of counts for a single cell, $z_{i}$ denotes the latent representation of the cell, and $\rho_{i}$ denotes the normalized expression of the cell. $f_{\theta}$ is formulated as a neural network with a final softmax layer. $h_{\phi}$ is a neural network used to parameterize zero-inflation probabilities for the generative zero-inflated negative binomial (ZINB) distribution. As such, for a single batch, the formulation of scVI is equivalent to the VAE with a ZINB likelihood. While all other models can be extended easily, DISCoVeR requires reconstructions as a proxy for the adversarial loss. For this formulation, we directly treat the normalized expressions $\rho_{i}$ as the adversarial reconstructions $\hat{x}$.