### v2

# Variational Learning of Disentangled Representations

Abstract
Disentangled representations allow models to separate factors shared across conditions from those that are condition-specific. This separation is crucial in domains such as biomedicine, where generalization to new treatments, patients, or species requires isolating stable biological signals from context-dependent effects. While several VAE-based extensions aim to achieve this, they often exhibit leakage between latent variables, limiting generalization. We introduce DisCoVR, a variational framework that explicitly separates invariant and condition-specific factors through: (i) a dual-latent architecture, (ii) parallel reconstructions to keep both representations informative, and (iii) a max–min objective that enforces separation without handcrafted priors. We show this objective maximizes data likelihood, promotes disentanglement, and admits a unique equilibrium. Empirically, DisCoVR achieves stronger disentanglement on synthetic data, natural images, and single-cell RNA-seq datasets, establishing it as a principled approach for multi-condition representation learning.
1 Introduction
Neural network–based models excel at learning rich representations of complex data and are increasingly applied in settings where each data point $x\in\mathcal{X}\subseteq\mathbb{R}^{d}$ is associated with a condition label $y\in{1,\dots,K}$. In biology, for example, conditions may represent treatments, patients, or species. Such representations are valuable for tasks like domain adaptation and transfer learning (transfer), where models must generalize from observed to novel conditions. Achieving this requires disentangled representations that separate factors shared across conditions from those specific to each $y$.
Generative models provide a natural framework for uncovering latent structure, with prominent examples including Generative Adversarial Networks (GANs) (goodfellow2020generative), Variational Autoencoders (VAEs) (vae), and diffusion models (sohl2015deep; ho2020denoising). Among these, VAEs and their extensions are particularly well-suited to transfer learning and domain adaptation (akrami2020brain; lovric2021should; godinez2022design; zhang2023variational), thanks to their probabilistic foundation and ability to capture uncertainty.
Thus, several VAE-based methods have been proposed to integrate data across multiple conditions or sources (xu2021probabilistic; lotfollahi2019scgen; boyeau2022deep), but only a few explicitly disentangle shared and condition-specific components (CVAE; CSVAE; CCVAE). While these approaches improve separation to some degree, they often depend on handcrafted priors that are difficult to design in high-dimensional domains like single-cell genomics, and they frequently exhibit residual leakage between latent varaibles, limiting generalization across conditions.
In this work, we introduce a framework for learning disentangled representations in multi-condition datasets. Our main contributions are: (i) a method combining two distinct reconstruction objectives with adversarial learning, reducing reliance on restrictive priors or handcrafted components; (ii) a max–min formulation of disentangled representation learning, along with a corresponding objective and theoretical guarantees for its equilibrium; and (iii) through experiments on synthetic benchmarks and real-world datasets, we show that our approach consistently improves upon existing methods disentanglement of shared and condition-specific structure.
2 DisCoVR: Disentangling Common and Variant Representations
For the task of learning disentangled representations from multi-condition data, we consider a dataset $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{n}$ consisting of inputs $x_{i}\in\mathcal{X}\subseteq\mathbb{R}^{d}$ collected from associated condition labels $y_{i}\in\{1,\dots,K\}$. For each class $k$ (corresponding to a study or experimental condition), the associated subset $\mathcal{D}_{k}\coloneqq\{x_{i}:y_{i}=k\}$ consists of i.i.d. samples drawn from a class-conditional distribution $p(x\mid y=k)$.
2.1 Model assumptions
We assume that the data is generated by latent variables $z$ and $w$, such that the joint distribution $p(x,y,z,w)$ factorizes according to the probabilistic graphical model illustrated in Figure 1(a), i.e.,
$p(x,y,z,w)=p(y)\,p(w\mid y)\,p(z)\,p(x\mid z,w).$ (1)
yywwzzxx
[…]
Figure 1: Graphical overview of our model. (a) Probabilistic graphical model: gray circles denote observed variables, white show latent variables. (b) Encoder–decoder architecture: the inhibition arrow from $\hat{y}$ to $\hat{x}$ corresponds to the adversarial component.
This model encodes two key conditional independence assumptions:
  1. 1.
    Latent variable conditional independence: Given the condition $y$, the latent representations $z$ and $w$ are conditionally independent: $z\perp w\mid y$.
  2. 2.
    Sufficiency of the shared latent representation: The input $x$ is conditionally independent of the condition $y$ given $w$: $x\perp y\mid w$.
However, note that in this formulation, $z$ and $w$ are no longer independent if conditioned also on $x$, that is, $z\not\perp w\mid x,y$.
2.2 Target posterior structure
In our model, each observation $x$ is generated from two latent variables: $z$, which is condition-invariant, and $w$, which is condition-aware through its dependence on $y$. Our goal is to learn probabilistic representations where the marginals of $z$ and $w$ preserve this structure, yielding disentangled factors. Thus, we seek to approximate the posterior $p_{z,w\mid x,y}$.
However, approximating the full posterior with a variational distribution $q_{z,w|x,y}$ is intractable: even for simple variational families such as Gaussians, modeling the dependencies between $z$ and $w$ requires a full covariance structure, which is computationally prohibitive in high dimensions. To mitigate this, we employ a factorized approximation $q_{z|x}\ q_{w|x,y}$.
Our variational approximation is guided by two complementary objectives: (i) $q_{z|x}$ closely approximating the marginal posterior $p_{z|x}$; and (ii) the product $q_{z|x}\ q_{w|x,y}$ closely approximating the true posterior $p_{z,w|x,y}$. Formally, given variational families [^1] Here we consider general families and specify our concrete choices in §2.4. $\mathcal{Q}_{z}$ and $\mathcal{Q}_{w}$ we seek to find $q^{*}_{z|x}\in\mathcal{Q}_{z}$ and $q^{*}_{w|x,y}\in\mathcal{Q}_{w}$ that minimize the following sum of Kullback-Leibler (KL) divergences:
$q^{*}_{z|x},\ q^{*}_{w|x,y}=\arg\min_{\begin{subarray}{c}q_{z|x}\in\mathcal{Q}_{z}\\ q_{w|x,y}\in\mathcal{Q}_{w}\end{subarray}}D_{\mathrm{KL}}(q_{z|x}\ \|\ p_{z|x})+D_{\mathrm{KL}}(q_{z|x}q_{w|x,y}\ \|\ p_{z,w|x,y}).$ (2)
2.3 Optimization objective
Since direct evaluation of the KL divergences in Equation 2 is intractable, we optimize a surrogate objective consisting of two ELBO terms.
The corresponding ELBO objective to minimize $D_{\mathrm{KL}}(q_{z|x}\ \|\ p_{z|x})$ is
$\mathcal{L}_{z}(q_{z|x},p;x)\coloneqq\mathbb{E}_{q_{z|x}}\left[\log p(x\mid z)\right]-D_{\mathrm{KL}}(q_{z|x}\,\|\,p_{z}),$ (3)
and the ELBO objective for the second KL term, $D_{\mathrm{KL}}(q_{z|x}q_{w|x,y}\ \|\ p_{z,w|x,y})$, is given by
$\textstyle\mathcal{L}_{w}(q_{w|x,y},p;x,y)\!\coloneqq$ $\mathbb{E}_{q_{z|x}}\!\left[\mathbb{E}_{q_{w|x,y}}\!\left[\log\left(p(x|z,w)\right)\right]\right]\!\!-\!D_{\mathrm{KL}}\left(q_{z|x}\,\|\,p_{z}\right)\!\!-\!D_{\mathrm{KL}}\left(q_{w|x,y}\,\|\,p_{w\mid y}\right).$ (4)
Note that $\mathcal{L}_{w}(q_{w|x,y},p;x,y)$ is the ELBO objective that corresponds to a factorized posterior $q_{z|x}q_{w|x,y}$. In Proposition 2.1 we examine the gap between this objective, and an ELBO term corresponding to a full variational posterior $q_{z,w|x,y}$. This can be interpreted as the cost of enforcing a condition-invariant latent representation, specifically, constraining $z$ to depend only on $x$. Proposition
Proposition 2.1.
For random variables $x,y,z,w$ following the graphical model in Figure 1(a),
$\operatorname{ELBO}(q,p;x,y)-\mathcal{L}_{w}(q_{w|x,y},p;x,y)=\mathbb{E}_{q_{w|x,y}}\left[\text{KL}\left(q_{z|x}\,\|\,p_{z|w,x,y}\right)\right].$
where $\operatorname{ELBO}(q,p;x,y)\coloneqq\log p(x\mid y)-D_{\mathrm{KL}}\left(q_{w|x,y}\,\|\,p_{w|x,y}\right)$.
The proof is provided in Appendix B.1.
Note that a full definition of the objectives in Equations 3 and 4 requires the specification of corresponding prior distributions, namely $p_{z}$ and $p_{w|y}$. We defer their definitions to §2.4.
Equation 4 provides an evidence lower bound on the conditional log-likelihood $\log p(x\mid y)$. By adding $\log p(y)$, this bound extends to the joint log-marginal likelihood $\log p(x,y)$. Beyond optimizing this objective, we aim to ensure that the marginal distribution over $y$ implicitly induced by the latent representations is consistent with the true $p(y)$.
To this end, we introduce an auxiliary classifier $g(y\mid z)$ as a form of posterior regularization (ganchev2010posterior). This classifier captures the residual predictive signal about $y$ in $z$ and is trained by minimizing the expected negative log-likelihood $\mathbb{E}_{q(z\mid x)}\log g(y\mid z)$. If $z$ is truly independent of $y$, then $g(y\mid z)$ will approximate the marginal distribution $p(y)$. By penalizing deviations from this behavior, we enforce the structural constraint $z\perp y$ in the learned representation.
Note that for this term to effectively encourage $q_{z|x}$ to discard condition-specific information, the classifier $g_{y|z}\in\mathcal{G}$ must be trained adversarially, with its own update step. This prevents degenerate solutions in which the loss is minimized without removing information about $y$ from $z$, for example, by collapsing $g$ to a constant predictor that ignores its input.
Combining the three terms above, we define the objective
$\mathcal{L}(q_{z|x},q_{w|x,y},g_{y|z};x,y)=\mathcal{L}_{z}(q_{z|x},p;x)+\mathcal{L}_{w}(q_{w|x,y},p;x,y)-\mathbb{E}_{q_{z|x}}\log g(y\mid z),$ (5)
which can be explicitly expressed as
$\mathcal{L}(q_{z|x},q_{w|x,y},g_{y|z};x,y)\coloneqq$ $\mathbb{E}_{q_{z|x}}[\log p(x\mid z)]+\mathbb{E}_{q_{z|x}}\left[\mathbb{E}_{q_{w|x,y}}[\log p(x\mid z,w)]\right]$
$-\mathbb{E}_{q_{z|x}}[\log g(y\mid z)]-2D_{\mathrm{KL}}(q_{z|x}\,\|\,p_{z})-D_{\mathrm{KL}}\left(q_{w|x,y}\,\|\,p_{w\mid y}\right).$ (6)
Finally, to enable flexible trade-offs between reconstruction expressiveness and disentanglement, we introduce weighting terms $\alpha=(\alpha_{1},\alpha_{2})$ into the objective following the motivation of $\beta$-VAEs (beta_vae):
$\mathcal{L}_{\alpha}(q_{z|x},q_{w|x,y},g_{y|z};x,y)\coloneqq$ $\mathbb{E}_{q_{z|x}}[\log p(x\mid z)]+\mathbb{E}_{q_{z|x}}\left[\mathbb{E}_{q_{w|x,y}}[\log p(x\mid z,w)]\right]$ (7)
$-\mathbb{E}_{q_{z|x}}\log g(y\mid z)-\alpha_{1}D_{\mathrm{KL}}(q_{z|x}\,\|\,p_{z})-\alpha_{2}D_{\mathrm{KL}}\left(q_{w|x,y}\,\|\,p_{w\mid y}\right).$
Accordingly, the mean weighted objective is suitable for max-min optimization of the form:
$\max_{q_{z|x}\in\mathcal{Q}_{z}}\max_{q_{w|x,y}\in\mathcal{Q}_{w}}\min_{g_{y|z}\in\mathcal{G}}\mathbb{E}_{p_{x,y}}\left[\mathcal{L}_{\alpha}(q_{z|x},q_{w|x,y},g_{y|z};x,y)\right].$ (8)
2.4 Latent prior models and variational approximations
Prior specification
We place a standard Normal prior over the condition-invariant latent variable, $p_{z}=\mathcal{N}(0,I)$, which reflects a non-informative prior belief over its values.
For the condition-aware latent variable $w$, we define a class-conditional Gaussian prior $p_{w|y}$. As a simple choice, we let $w$ have the same dimensionality as $z$ and specify
$p(w\mid y=k)=\mathcal{N}(\mu_{k},I),\qquad\mu_{k}\coloneqq\mathbb{E}_{p_{x|y=k}}\left[\mathbb{E}_{q_{z|x}}[z]\right].$ (9)
Here $\mu_{k}$ is the mean of the inferred latent representations $z$ within the $k$-th class [^2] Similarly, for different dimensions of $z$ and $w$ the mean aggregation can be replaced with a neural network that maps the inferred representations $z$ for each class to the parameters of the Gaussian prior..
This specification induces a coupling between the two latent variables through the data distribution. Aligning $p_{w|y}$ with the class-wise expectations of the invariant variable, further encourages $q_{z|x}$ to encode informative representations, since capturing the shared structures will now not only increase $\mathcal{L}_{z}(q_{z|x},p;x)$, but also decrease $D_{\mathrm{KL}}\left(q_{w\mid x,y}\,\|\,p_{w|y}\right)$, and thus increase $\mathcal{L}_{w}(q_{w|x,y},p;x,y)$.
Importantly, for a truly condition-agnostic $q_{z|x}$, the conditional expectations $\mu_{k}$ will collapse to a shared mean $\mu\coloneqq\mathbb{E}_{p_{x}}\left[\mathbb{E}_{q_{z|x}}\left[z\right]\right]$. In this case $p_{w|y}$ becomes a shared prior across classes, centered at a meaningful point in the latent space, rather than an uninformative one.
As the following proposition establishes, this anchoring of the prior $p_{w|y}$ in the variational distribution $q_{z|x}$ preserves the convex–concave structure of the objective, ensuring that the resulting max-min problem has a unique optimal solution.
Proposition 2.2.
Let $\mathcal{Q}_{z}$ and $\mathcal{Q}_{w}$ be convex parametric families of variational distributions over $z$ and $w$, respectively, and let $\mathcal{G}$ denote a convex set of classifiers such that $g(x)\in[0,1]$ for all $g\in\mathcal{G}$. Assume the latent priors are given by $z\sim p(z)$ and $p(w|y)=\mathcal{N}(\mu_{y},I)$, where $p(z)$ is a continuous strictly positive distribution, and $\mu_{y}=\mathbb{E}_{p_{x|y}}\left[\mathbb{E}_{q_{z|x}}[z]\right]$. Then, under standard regularity conditions (see Appendix B.2.1), there exists a unique saddle point:
$\left(q_{z|x}^{*},q_{w|x,y}^{*},g^{*}_{y\mid z}\right)=\max_{q_{z|x}\in\mathcal{Q}_{z}}\max_{q_{w|x,y}\in\mathcal{Q}_{w}}\min_{g_{y|z}\in\mathcal{G}}\mathcal{L}(q_{z|x},q_{w|x,y},g_{y|z}).$
The proof is provided in Appendix B.2.2.
Specification of variational families
We set both variational families $\mathcal{Q}_{z}$ and $\mathcal{Q}_{w}$ as $d$-dimensional Gaussian distributions with diagonal covariance matrices. Accordingly, each variational distribution is parameterized by a mean vector $\mu\in\mathbb{R}^{d}$ and a vector of variances $\sigma^{2}\in\mathbb{R}^{d}_{+}$ corresponding to the diagonal of the covariance matrix, yielding $\theta=(\mu,\sigma^{2})$.
3 Encoder-decoder model
In order to optimize the objective in Equation 8 with respect to $q_{z|x},\,q_{w|x,y}$, and $g_{y|z}$ over the dataset $\mathcal{D}$, we introduce an encoder-decoder framework (illustrated in Figure 1(b)). In this framework, two separate reconstructions of $x$ are generated: one, denoted $\hat{x}\sim p_{x|z}$, where $z$ is sampled from the condition-invariant posterior $q_{z|x}$, and the other, denoted $\tilde{x}\sim p_{x|z,w}$, where in addition, $w$ is sampled from the condition-aware posterior $q_{w|x,y}$. The corresponding algorithm is summarized in Algorithm 1.
Condition agnostic representation   An input $x\in\mathcal{X}$ is mapped to the variational parameters $\theta_{z}=(\mu_{z},\sigma^{2}_{z})$ by an encoder neural network $f^{z}_{\phi}:\mathcal{X}\rightarrow\mathbb{R}^{d}\times\mathbb{R}^{d}_{+}$ parametrized by weights $\phi$. A latent encoding $z\sim q_{\theta_{z}}$ is then sampled and mapped to a reconstruction $\hat{x}$ via a decoder neural network $h^{z}_{\psi}:\mathbb{R}^{d}\rightarrow\mathcal{X}$ parametrized by weights $\psi$.
Adversarial classifier   Rather than learning a complex classifier from $z$ to $y$, we use the reconstruction $\hat{x}$ from $z$, and predict $y$ from $\hat{x}$ via a simple multinomial logistic regression $g_{\beta}:\hat{\mathcal{X}}\rightarrow[0,1]^{K}$, with class-specific weights $\beta={\beta_{k}}_{k=1}^{K}$, or a shallow MLP. Since $\hat{x}$ is a deterministic function of $z$, this is equivalent to applying a restricted classifier on $z$. By the data processing inequality, such a classifier can only capture a subset of the information $z$ contains about $y$; thus, maximizing this lower bound on $I(z;y)$ also maximizes $I(z;y)$ itself. Although this substitution weakens the estimation of the cross-entropy term $\mathbb{E}{q{z\mid x}}\log g(y\mid z)$, from an information-theoretic standpoint, we observed this to be often advantageous in practice: predicting $y$ from $\hat{x}$ reduces the variance introduced by sampling $z\sim q_{\theta_{z}}$, providing a regularizing effect that prevents $q_{\theta_{z}}$ from overfitting to noisy classifier signals.
Condition aware representation   A labeled input pair $(x,y)\in\mathcal{X}\times\{1,\dots,K\}$ is mapped to the parameters $\theta_{w}=(\mu_{w},\sigma^{2}_{w})$ using an encoder neural-network $f^{w}_{\rho}:\mathcal{X}\times\{1,\dots,K\}\rightarrow\mathbb{R}^{d}\times\mathbb{R}^{d}_{+}$ parametrized by weights $\rho$. A sample $w\sim q_{\theta_{w}}$ is then drawn, and the pair $(z,w)$ is mapped to a reconstruction $\tilde{x}$ via a decoder neural-network $h^{z,w}_{\eta}:\mathbb{R}^{d}\rightarrow\mathcal{X}$, parametrized by weights $\eta$. To compute the prior $p_{w|y}$, we estimate the class-specific mean as $\hat{\mu}_{k}\coloneqq\frac{1}{n_{k}}\sum_{i;y_{i}=k}z_{i}$ where each $z_{i}\sim p(z\mid x_{i})$ is sampled from the encoder given an input $x_{i}$ with label $y_{i}=k$, and $n_{k}$ is the number of training points with the label $y=k$.
Algorithm 1
1:Input: Data $\mathcal{D}=\{x_{1:n},y_{1:n}\}$, number of training iterations $J$, initial parameters $\phi^{(0)},\psi^{(0)},\rho^{(0)},\eta^{(0)},\beta^{(0)}$, learning rates $\gamma_{1},\gamma_{2}$, weighting terms $\alpha=(\alpha_{1},\alpha_{2})$
2:for $1\leq j\leq J$ do
3:  Compute $\theta_{z}=f^{z}_{\phi^{z}}(x)$ and $\theta_{w}=f^{w}_{\rho^{(j-1)}}(x,y)$
4:  Sample condition invariant and aware latent variables $z\sim q_{\theta_{z}}$ and $w\sim w_{\theta_{w}}$
5:  Compute reconstructions $\hat{x}=h^{z}_{\psi^{(j-1)}}(z)$ and $\tilde{x}=h^{z,w}_{\eta^{(j-1)}}$
[…]
7:  Update classifier parameters:
$\beta^{(j)}\leftarrow\beta^{(j-1)}-\gamma_{1}\nabla_{\beta}\,\mathcal{L}_{\alpha}(q_{z|x},q_{w|x,y},g_{y|z})\Big|_{\phi=\phi^{(j-1)},\psi=\psi^{(j-1)},\rho=\rho^{(j-1)},\eta=\eta^{(j-1)}}$
8:  Update encoder and decoder parameters, $\Omega^{(j)}=\left(\phi^{(j)},\psi^{(j)},\rho^{(j)},\eta^{(j)}\right),$
$\Omega^{(j)}\leftarrow\Omega^{(j-1)}-\gamma_{2}\nabla_{\phi,\psi,\rho,\eta}\,\mathcal{L}_{\alpha}(q_{z|x},q_{w|x,y},g_{y|z})\Big|_{\beta=\beta^{(j)}}$
9:end forReturn: $\phi^{(J)},\psi^{(J)},\rho^{(J)},\eta^{(J)},\beta^{(J)}$
In practice, following standard approaches in VAE-based models, (i) we use a single-sample Monte Carlo estimate to approximate the expectations in Equation 8, and (ii) instead of directly sampling from $q_{\theta}$, we employ the reparameterization trick to enable differentiable sampling. Specifically, we sample $\epsilon\sim\mathcal{N}(0,I)$ and obtain a sample from $q_{\theta}$ by applying a deterministic transformation of $\epsilon$ based on the variational parameters $\theta$.
4 Comparison to previous approaches
Here we review VAE-based methods for disentangled representation learning, which form the primary basis for comparison with our approach. Broader related literature is discussed in Appendix A.
VAEs (vae) are generative models that learn latent representations by maximizing the evidence lower bound (ELBO) on the data log-likelihood:
$\mathbb{E}_{q_{z|x}}[\log p(x\mid z)]-D_{\mathrm{KL}}\left(q_{z\mid x}\ \|\ p_{z}\right)\leq\log p(x),$
where $(x,z)\sim p$, and $z|x\sim q$ is a latent variable inferred from data. VAEs consist of an encoder $q_{z\mid x}$ that maps inputs to latent distributions, and a decoder $p_{x\mid z}$ that reconstructs inputs from latent representations. The learning process frames posterior inference as KL-regularized optimization over a variational family $\mathcal{Q}$ aiming to approximate the posterior $p_{z|x}$ under a typically simple prior $p(z)$. Several VAE extensions were proposed to encourage disentanglement. These are discussed below.
Conditional VAEs   (CVAE) incorporate supervision to the standard VAE model by conditioning both the encoder and decoder on an observed label $y$, yielding the following objective:
$\mathbb{E}_{q_{z|x,y}}[\log p(x\mid z,y)]-D_{\mathrm{KL}}\left(q(z\mid x,y)\|p(z)\right)\leq\log p(x|y).$
While this allows controlled generation and partial disentanglement between $z$ and $y$, the latent variable $z$ is still inferred from both $x$ and $y$ and thus can encode condition-specific information.
Conditional Subspace VAEs  (CSVAEs) (CSVAE), explicitly factorize the latent space into a shared component $z$ and a label-specific component $w$ (see Supplementary Figure 1(a)). Similarly, their hierarchical extension beker2024patches introduces an intermediate latent variable between $x$ and $(z,w)$. As in our approach, to encourage disentanglement, CSVAEs introduce an adversarial regularization term that penalizes mutual information between $z$ and $y$, thereby discouraging predictability of $y$ from $z$. They are learned by optimizing the following lower bound on $\log p(x\mid y)$:
$\textstyle\mathbb{E}_{q_{z,w|x,y}}[\log p(x\mid w,y)]\!-\!\mathbb{E}_{q_{z|x}}[\int g(y\mid z)\log g(y\mid z)\,dy]\!-\!D_{\mathrm{KL}}\left(q_{w\mid x,y}\,\|\,q_{w\mid y}\right)\!-\!D_{\mathrm{KL}}\left(q_{z\mid x}\,\|\,p_{z}\right).$
Domain Invariant VAEs (DIVA) and Characteristic-capturing VAEs (CCVAE)   DIVA (diva) and CCVAE (CCVAE), shown in Supplementary Figure 1(b), introduce two latent variables, $z$ and $w$, where $w$ captures label-related features by jointly optimizing a classifier $p(w\mid y)$ jointly alongside the remaining objective. For fully supervised cases, the DIVA model optimizes
$\textstyle\mathbb{E}_{q_{z,w|x}}[\log p(x\mid z,w)]+\mathbb{E}_{q_{w|x}}[\log q(y\mid w)]-D_{\mathrm{KL}}\left(q_{z|x}\|p_{z}\right)-D_{\mathrm{KL}}\left(q_{w\mid x}\|p_{w\mid y}\right)\leq\log p(x,y).$
This objective corresponds to the assumptions $x\perp y\mid z,w$ and $z\perp w$. Similarly, CCVAEs optimize the following lower bound on $\log p(x\mid y)$:
$\textstyle\mathbb{E}_{q_{z,w|x,y}}\left[\frac{q(y\mid w)}{q(y\mid x)}\log\frac{p(x\mid z,w)}{q(y\mid w)}\right]-D_{\mathrm{KL}}\left(q_{z|x}\|p_{z|y}\right)-D_{\mathrm{KL}}\left(q_{w\mid x}\|p_{w\mid y}\right)+\log q(y\mid x).$
Comparison:   In prior methods, reconstruction is performed jointly from both representations $z$ and $w$, via $p(x\mid z,w)$. This design provides no incentive to distribute information meaningfully between $z$ and $w$: the model can place all relevant information into $w$, leaving $z$ either uninformative or entangled with $w$. Our method addresses this limitation through two key components: (i) a separate reconstruction term from $z$ alone, which explicitly forces $z$ to capture informative, condition-invariant structure. Theoretically, we show that the additional term $p(\hat{x}\mid z)$ is required to bound the gap between our approximate posterior and the full model ELBO under the factorization assumption; (ii) a prior over $w$ conditioned on the mean of $z$, which, through label-specific aggregates, discourages leakage of class-invariant information into $w$ and reduces redundancy between the two representations.
Another novelty of our approach is a principled probabilistic objective that enforces the correct conditional independence. Without an explicit probabilistic model, prior methods sought to enforce $z\perp w\mid x,y$. In contrast, our formulation shows that the proper requirement is the weaker condition $z\perp w\mid y$. As established in Section 2.1, our model satisfies $z\perp w\mid y$, but not $z\perp w\mid x,y$.
To enforce this criterion, we propose a theoretically grounded variational objective that uses a conditional ELBO for the dependent representation $w$, with the prior over $w$ conditioned on the mean of the independent representation $z$.
5 Experiments
Datasets: We evaluate DisCoVR against existing approaches on synthetic data, natural images, and biological data. These datasets were chosen to probe condition-invariant structure and to ensure comparability with prior work: for instance, Swiss rolls and CelebA were used in CSVAE, and CelebA also in CCVAE.
Evaluation: When applicable, we evaluate reconstruction quality using negative log-likelihood (NLL), root mean squared error (RMSE), and the absolute deviation from the optimal-Bayes classifier on the reconstructed data, denoted as $\Delta$–Bayes. Disentanglement is quantified via a neural estimator of the mutual information $I(z;w)$ (pmlr-v80-belghazi18a). Full model architectures, hyperparameters, and additional implementation details are provided in Appendix H. Our results show that DisCoVR achieves superior performance across all experiments.
5.1 Simulated data
We begin with controlled synthetic experiments to isolate and visualize disentanglement.
5.1.1 Parametric model
Data generating model: Consider a model where the observed data $x$ is generated as a function of two latent variables $z$ and $w$, and $y$ are binary labels. Assume that the marginal distributions of the latent variables are given by $z\sim\mathcal{N}(0,1)$ and $w\sim\mathcal{N}(0,1)$, and that the data $x$ is generated as the sum of the two latent variables: $x=z+w$. Since $z$ and $w$ are both drawn from $\mathcal{N}(0,1)$, it follows that $x\sim\mathcal{N}(0,2)$. Finally, assume that the binary label is determined by the sign of $w$: $y=1$ if $w>0$, and $y=0$ otherwise.
Optimal disentanglement: Given that $z$ and $w$ are independent and $x=z+w$, we have that $p(z\mid x)=\mathcal{N}\!\left(z;\frac{x}{2},\frac{1}{2}\right)$. Hence, given $x$, the best estimate for $z$ is $\frac{x}{2}$. Note that when ignoring the label $y$, the conditional distribution $p(w\mid x)$ is $p(w\mid x)=\mathcal{N}\!\left(w;\frac{x}{2},\frac{1}{2}\right)$. However, the observation of $y$ (which indicates whether $w$ is positive or negative) truncates this distribution:
$p(w\mid x,y=1)=\frac{\mathcal{N}\left(w;\frac{x}{2},\frac{1}{2}\right)}{\Phi\left(\frac{x}{\sqrt{2}}\right)},\qquad p(w\mid x,y=0)=\frac{\mathcal{N}\left(w;\frac{x}{2},\frac{1}{2}\right)}{1-\Phi\left(\frac{x}{\sqrt{2}}\right)}.$ (10)
Results: Table 1 shows that DisCoVR (ours) best approximates the analytic posteriors, resulting in the lowest deviation from the optimal Bayes classifier and the best reconstruction.
Table 1: Parametric model results: DisCoVR (ours) outperforms all competitors across all metrics.
NLL $\downarrow$ $D_{\mathrm{KL}}(q_{z\mid x}\;||\>p_{z\mid x})\downarrow$ $D_{\mathrm{KL}}(q_{w\mid x,y}\;||\>p_{w\mid x,y})\downarrow$ $\Delta$ – Bayes $\downarrow$
CSVAE - N.A. $1.810\pm 0.016$ $6.65\pm 3.46$ $23.61\pm 0.36$ $24.83\pm 0.04$
CSVAE $1.786\pm 0.022$ $2.85\pm 1.11$ $23.98\pm 4.36$ $24.33\pm 1.28$
HCSVAE - N.A. $1.784\pm 0.010$ $4.01\pm 0.07$ $25.82\pm 0.38$ $24.99\pm 0.01$
HCSVAE $1.770\pm 0.004$ $3.99\pm 0.09$ $26.25\pm 0.59$ $24.99\pm 0.01$
DIVA $1.788\pm 0.008$ $3.21\pm 1.52$ $12.88\pm 3.31$ $3.51\pm 0.32$
CCVAE $1.785\pm 0.006$ $1.77\pm 0.81$ $12.95\pm 3.35$ $3.57\pm 0.15$
DisCoVR (ours) $\mathbf{1.769\pm 0.003}$ $\mathbf{0.17\pm 0.01}$ $\mathbf{10.10\pm 0.73}$ $\mathbf{0.1\pm 0.28}$
5.1.2 Noisy Swiss roll
Dataset:   We use a noisy variant of the labeled Swiss Roll dataset (swissroll; CSVAE), generating $n=20,000$ and assigning binary labels based on a lengthwise split, with labels flipped at rate $\rho$. The common geometry (its projection along the 2D plane) remains intact, the conditional structure along the third axis becomes noisy. Figure 2A illustrates the setup.
Table 2: Noisy Swiss roll results: DisCoVR (ours) yields lowest deviation from optimal-Bayes, maintains low latent leakage, and high reconstruction accuracy.
$I(z;w)\downarrow$ NLL $\downarrow$ $\Delta$ – Bayes $\downarrow$
CSVAE - N.A. $0.047\pm 0.023$ $3.303\pm 0.003$ $23.88\pm 12.02$
CSVAE $0.031\pm 0.025$ $\mathbf{3.302\pm 0.003}$ $17.99\pm 14.58$
HCSVAE - N.A. $0.024\pm 0.012$ $\mathbf{3.302\pm 0.002}$ $30.00\pm 0.00$
HCSVAE $\mathbf{0.002\pm 0.001}$ $\mathbf{3.302\pm 0.002}$ $30.00\pm 0.00$
DIVA $0.336\pm 0.083$ $\mathbf{3.302\pm 0.003}$ $1.88\pm 1.05$
CCVAE $0.502\pm 0.089$ $\mathbf{3.302\pm 0.002}$ $2.21\pm 0.84$
DisCoVR (Ours) $0.005\pm 0.002$ $\mathbf{3.302\pm 0.002}$ $\mathbf{1.14\pm 0.21}$
Optimal disentanglement:   Since the Swiss Roll is sliced at the center and label noise is applied uniformly, marginalizing over labels yields a symmetric spiral centered along the roll—i.e., the marginal posterior $p(z\mid x)$} is label-invariant. In contrast, the conditional component retains a noisy but informative signal, with a uniform noise level of $\rho=0.3$. As a result, the Bayes optimal classifier trained on any realistic representation is upper-bounded at 70% accuracy. Results:   Figure 2 presents qualitative and quantitative results, showing that DisCoVR both models the label noise accurately and effectively disentangles shared and condition-specific structure. Notably, DisCoVR captures the marginal data distribution, successfully recovering the expected spiral pattern, as shown in Figure 2B (right). Additionally, the results in Table 2 show that DisCoVR achieves the lowest deviation from the optimal Bayes classifier and minimal information leakage between latent variables, while preserving reconstruction quality.
Figure 2: A: Noisy labeled Swiss Roll dataset. B (left): Reconstruction and conditional embedding from DisCoVR. B (right): Marginal reconstruction from the shared embedding, recovering the spiral structure. C: Comparison of reconstruction and conditional embeddings across models.
5.2 Real data
5.2.1 Noisy colored MNIST
Dataset:   We use a modified MNIST (deng2012mnist) dataset constructed from $60,000$ duplicated images: in one copy we remove the red channel ($y=0$) and in the other we remove the green channel ($y=1$). The digit shape remains intact following the blue channel. Label noise is introduced by flipping labels with probability $\rho\in\{0,0.1,0.2,0.3,0.4\}$.
Optimal disentanglement:   Since the colored images are generated in equal proportions, the marginal reconstruction for retains a single color (see Figure 3).
Results:   We evaluate marginal coloring reconstruction by DisCoVR and previous methods. Under no label noise ($\rho=0$), all methods perform similarly (Supplementary Figure 3). At all non-zero noise levels, DisCoVR consistently outperforms competitors, whose marginal reconstructions are averaged over class-conditioned outputs. Metrics for $\rho=0.3$ are shown in Supplementary Table 2, with results for other noise levels in Supplementary Figure 4.
Figure 3: Colored MNIST reconstructions from the label-agnostic representation $z$ for noise level $\rho=0.3$. DisCoVR is the only model that consistently reconstructs mixed “semi-red/blue” tone (purple) indicating that color information has been removed and matching the true marginal $p(x|B)$.
5.2.2 CelebA-Glasses
Dataset:   We use all CelebA (liu2015faceattributes) images labeled with eyeglasses attribute ($y=1$), and twice as many randomly sampled images without ($y=0$), totaling $n=35,712$ samples.
Results:   Figure 4 shows that DisCoVR accurately reconstructs input images while producing shared embeddings that marginalize over the eyeglasses attribute, consistently adding ”pseudo-glasses” to all samples. Competing methods are shown in Supplementary Figure 5, with quantitative results in Supplementary Table 3. While reconstruction quality is comparable across methods, DisCoVR achieves significantly better disentanglement.
Figure 4: CelebA-Eyeglasses results. Top: Original images with/without eyeglasses. Middle: Full reconstructions by DisCoVR. Bottom: Reconstructions solely from common embeddings $z$. A shared representation needs to be invariant to $y$ (presence or absence of glasses). Indeed, all reconstructed faces display an intermediate ”semi-glasses” appearance, regardless of the original label.
Results for an additional experiment with the wearing-hat attribute are provided in Appendix G.
5.2.3 Single cell RNA-sequencing (scRNA-seq) data from Lupus dataset
Dataset:   We analyze single-cell RNA sequencing from $n=13,999$ peripheral blood mononuclear cells (PBMCs) collected from 8 lupus patients under two conditions: 7,451 cells control ($y=0$), and 6,548 IFN-$\beta$ stimulation cells. IFN-$\beta$ stimulation induces notable shifts in gene expression, visible in the UMAP embedding in Figure 5B (left).
Figure 5: A–B (left): UMAPs of raw gene counts from the IFN-$\beta$ dataset. A–B (middle): Shared embedding $z$ aligns cells by type while removing stimulation effects. A–B (right): Condition-specific embedding $w$ isolates the stimulation effect. C: Reconstructions from both $z$ and $w$ (yellow) recover empirical gene means and variances, while reconstructions from $z$ alone (blue) miss the stimulation-induced variance, confirming that $z$ discards $y$ while preserving cell-type features.
Results:   Supplementary Table 17 shows that DisCoVR effectively achieves the desired behavior with strong empirical performance, where only cell type information is captured in $z$ (Figure 5A, middle) while the effects of IFN-$\beta$ stimulation are wholly represented in $w$ (Figure 5B, right). Other approaches either (1) achieve mixing in the $z$ space, but compromise on keeping cell types separated or (2) leak information about stimulation into the $z$ space (Supplementary Figure 6).
Facilitating interpretability:   By enabling marginalized reconstructions, DisCoVR provides a direct link between shared embeddings and gene expression, offering clearer insight into the effects of IFN-$\beta$ stimulation, unlike other methods. In Figure 5C, comparing variance across marginal and full reconstructions accurately recovers gene-level differences associated with IFN-$\beta$ stimulation, including ISG15, FTL, CCL8, CXCL10, CXCL11, APOBEC3A, IL1RN, IFITM3 and RSAD2.
6 Conclusion
In this work we introduced a variational framework for disentangled representation learning. Our formulation explicitly separates condition-invariant and condition-aware factors. Unlike prior work, DisCoVR incorporates two reconstruction paths, one based solely on the shared latent variable $z$, and the other on both latent variables $z$ and $w$. Our model simultaneously learns informative shared representations, and captures structured variation across conditions. Experimental results demonstrate that DisCoVR achieves strong reconstruction, minimal information leakage, and accurate modeling of conditional effects, consistently outperforming existing methods.
Appendix A Additional related work
A.1 Domain generalization
The task of representation disentanglement is closely related to the field of domain generalization (muandet2013domain), which assumes limited or no access to target domain samples and aims to learn representations that can be readily adapted, often via transfer learning, to new, unseen domains.
As noted by wang2019learning, existing methods in domain generalization can be broadly categorized into two main approaches: (i) approaches for reducing the inter-domain differences, often by using adversarial techniques (ghifary2015domain; wang2017select; motiian2017unified; li2018domain; carlucci2019; li2018domain; wang2019learning; akuzawa2020adversarial; zhu2022localized; gokhale2023improving; dayal2023madg; cheng2023adversarial; chen2024domain), and (ii) Approaches that construct an ensemble of domain-specific models, and then fuse their representations to form a unified, domain-agnostic representation (ding2017deep; mancini2018best; zhou2021domain; muhammad2024domain).
Additional strategies for domain generalization include contrastive learning approaches (kim2021selfreg), methods based on distribution alignment via metrics (muandet2013domain; sun2016deep), and techniques utilizing custom network architectures, for instance by incorporating domain-specific adapters between shared layers (rebuffi2017learning; rebuffi2018efficient; li2019efficient; omi2022model).
The primary distinction between these methods and ours lies in the explicit probabilistic modeling and disentanglement of domain-invariant and domain-specific factors. Whereas prior approaches typically focus on aligning domains through adversarial training or fusing multiple domain-specific predictors, our method constructs a structured latent space, decomposed into a shared representation $z$, capturing domain-invariant information, and a conditional component $w$, which encodes domain-specific variability. This factorization is learned through a tailored variational objective involving an adversarial penalty and two reconstructions —one based on $z$ alone, and another on the full latent pair $(z,w)$, thereby promoting both interpretability and a clean separation of shared and domain-aware features.
A.2 Out of distribution generalization
A.2.1 Enviroment balancing methods
The field of out-of-distribution (OOD) generalization emerged from foundational work on causality and invariance across training environments (peters2016causal; peters2017elements). The central assumption is that each environment exhibits distinct spurious correlations between features and labels; therefore, robust generalization requires models to focus on invariant relationships that hold across environments. To address this distribution shift, many recent approaches adopt a regularized empirical risk minimization framework:
$\min_{\theta}\quad\smashoperator[]{\sum_{e\in E_{\text{train}}}^{}}\ell^{e}(f_{\theta})+\lambda R(f_{\theta},E_{\text{train}}),$ (11)
where the regularizer $R$ encourages representations that are stable across environments. Among these, Invariant Risk Minimization (IRM) (arjovsky2019invariant) enforces that a single classifier remains optimal across all environments, Variance Risk Extrapolation (VarREx) (krueger2021out) promotes robustness by minimizing the variance of losses across environments, and CLOvE (wald2021calibration) takes a calibration-theoretic perspective, penalizing discrepancies between predicted confidence and correctness across environments.
While these methods focus on enforcing predictive invariance across environments through regularization, our approach instead explicitly enforces conditional independence between the shared latent variable $z$ and an environment-aware variable $w$.
A.2.2 Distributionally robust methods
An alternative line of work for handling distribution shifts is Distributionally Robust Optimization (DRO) (DRO; duchi2021statistics; duchi2021learning; wei2023distributionally), which avoids assuming a fixed data-generating distribution. Instead, DRO methods optimize performance under the worst-case scenario over a family of plausible distributions. A prominent variant, known as group DRO (sagawa2019distributionally; piratla2021focus), introduces group-level structure that may correlate with spurious features, potentially leading to biased predictions. In settings where group labels are not directly observed, several strategies have been proposed, including reweighting high-loss examples (liu2021just) and balancing class-group combinations through data sub-sampling (idrissi2022simple).
However, these approaches assume that the label space remains fixed between training and test time, limiting their applicability in adaptation to new domains, environments or conditions.
A.3 Zero-shot learning
Zero-shot learning systems (few_shot; zero_shot) aim to classify instances from novel, previously unseen classes at test time. In contrast to the out-of-distribution (OOD) generalization setting, these approaches typically do not assume the presence or structure of a distribution shift. Instead, a common strategy is to learn data representations that capture class-agnostic similarity, enabling the model to determine whether two instances belong to the same class without requiring knowledge of the class identity itself. Such methods include contrastive-learning (contrastive), siamese neural networks (siamese_networks), triplet networks (triplet_networks), and other more recent variations (oh2016deep; sohn2016improved; wu2017sampling; yuan2019signal). Recent work has begun to address the impact of class distribution shifts in zero-shot settings. For instance, slavutsky2024class integrate environment-based regularization—motivated by OOD generalization—with zero-shot learning by simulating distribution shifts through hierarchical sampling, enabling the model to learn representations that are robust to shifts in class distributions.
While this line of work shares our motivation of improving robustness under unseen conditions, it primarily addresses the problem of class-level generalization through similarity-based learning, rather than explicitly modeling and disentangling the latent factors—such as domain or environment—that drive distributional variation across tasks.
Appendix B Proofs
B.1 Proof of Proposition 2.1
Proof.
$\operatorname{ELBO}(q,p;x,y)-\mathcal{L}_{w}(q_{w|x,y},p;x,y)$ (12)
$=\left[\log p(x\mid y)-D_{\mathrm{KL}}(q_{w|x,y}\,\|\,p_{w|x,y})\right]-\left[\log p(x\mid y)-D_{\mathrm{KL}}\left(q_{z|x}q_{w|x,y}\,\|\,p_{z,w|x,y}\right)\right]$ (13)
$=D_{\mathrm{KL}}\left(q_{z|x}q_{w|x,y}\,\|\,p_{z,w|x,y}\right)-D_{\mathrm{KL}}\left(q_{w|x,y}\,\|\,p_{w|x,y}\right)$ (14)
$=\mathbb{E}_{q_{w|x,y}}\left[\mathbb{E}_{q_{z|x}}\left[\log q(z\mid x)+\log q(w|x,y)-\log p(z,w\mid x,y)\right]\right]$ (15)
$\phantom{=}-\mathbb{E}_{q_{w|x,y}}\left[\log q(w|x,y)-\log p(w|x,y)\right]$ (16)
$=\mathbb{E}_{q_{w|x,y}}\left[\mathbb{E}_{q_{z|x}}\left[\log q(z\mid x)-\log p(z,w\mid x,y)+\log p(w|x,y)\right]\right]$ (17)
$=\mathbb{E}_{q_{w|x,y}}\left[\mathbb{E}_{q_{z|x}}\left[\log q(z\mid x)-\log p(z\mid w\mid x,y)\right]\right]$ (18)
$=\mathbb{E}_{q_{w|x,y}}\left[\text{KL}\left(q_{z|x}\,\|\,p_{z|w,x,y}\right)\right].$ (19)
B.2 Game equilibrium
B.2.1 Regularity conditions
To ensure that expectations and KL‑terms in the game objective $\mathcal{L}(q_{z|x},q_{w|x,y},g_{y|z})$ render the functionals strictly concave in $q_{z|x}$ , strictly concave in $q_{w|x,y}$, and strictly convex in $g$, the following regularity conditions are required:
  1. 1.
    The likelihoods $p(x|z),p(x|z,w),p(y|x)$ are strictly positive, continuous densities.
  2. 2.
    The variational families $Q_{z}$ and $Q_{w}$, and the set of achievable classifiers $\mathcal{G}$ are non‑empty, convex and compact.
  3. 3.
    $\log p(x|z,w)$ and $\log g(y|x)$ are integrable.
B.2.2 Proof of Proposition 2.2
Proof.
Since $\mathcal{L}_{z}(q_{z|x},p;x)$ is the standard ELBO objective, we have that
$\mathcal{L}_{z}(q_{z|x},p;x)=\log p(x)-D_{\mathrm{KL}}\left(q_{z|x}\,\|\,p_{z|x}\right).$ (20)
Similarly, we have that
$\mathcal{L}_{w}(q_{w|x,y},p;x,y)$ $=\log p(x\mid y)-D_{\mathrm{KL}}\left(q_{z|x}q_{w|x,y}\,\|\,p_{z,w|x,y}\right).$ (21)
Thus,
$\mathcal{L}(q_{z|x},q_{w|x,y},g_{y|z})=\mathbb{E}_{p_{x,y}}$ $\left[\log p(x)-D_{\mathrm{KL}}\left(q_{z|x}\,\|\,p_{z|x}\right)\right$. (22)
$+\log p(x\mid y)-D_{\mathrm{KL}}\left(q_{z|x}q_{w|x,y}\,\|\,p_{z,w|x,y}\right)$ (23)
$\left.-\mathbb{E}_{q_{z|x}}\log g(y\mid z)\right].$ (24)
For fixed $q_{z|x}$, the adversarial classifier minimizes:
$\mathbb{E}_{p_{x,y}}\mathbb{E}_{q_{z|x}}\log g(y\mid z),$ (25)
which is the population cross-entropy and is strictly convex in $g(y|z)$, and thus has a unique solution.
It remains to show that the terms in the objective function that depend on $q_{z|x}$ and $q_{w|x,y}$, are strictly concave in each argument when the others are held fixed.
Focusing on the terms dependent on $q_{w|x,y}$ first, define
$\ell_{w}\coloneqq$ $-D_{\mathrm{KL}}\left(q_{z|x}q_{w|x,y}\,\|\,p_{z,w|x,y}\right)$ (26)
$=$ $-\iint q\left(z\mid x\right)q(w\mid x,y)\left[\log q\left(z\mid x\right)+\log q(w\mid x,y)-\log p(z,w\mid x,y)\right]\,dz\,dw$
$=$ $-\int q\left(z\mid x\right)\log q\left(z\mid x\right)\,dz+\int q(w\mid x,y)\log q(w\mid x,y)\,dw$ (27)
$-\iint q\left(z\mid x\right)q(w\mid x,y)\log p(z,w\mid x,y)\,dz\,dw$ (28)
$=$ $H(q_{z\mid x})+H(q_{w\mid x,y})+\mathbb{E}_{q_{z|x}}\mathbb{E}_{q_{w|x,y}}\log p(z,w\mid x,y).$ (29)
Note that
$\mathbb{E}_{q_{z|x}}\mathbb{E}_{q_{w|x,y}}\log p(z,w\mid x,y)$ (30)
is linear in $q_{w|x,y}$, and since $H(q_{w\mid x,y})$ is strictly concave in $q_{w|x,y}$, we have that $\mathbb{E}_{p_{x,y}}[\ell_{w}]$ is strictly concave in $q_{w|x,y}$.
Similarly, define
$\ell_{z}$ $\coloneqq-D_{\mathrm{KL}}\left(q_{z|x}\,\|\,p_{z|x}\right)-D_{\mathrm{KL}}\left(q_{z|x}q_{w|x,y}\,\|\,p_{z,w|x,y}\right).$ (31)
By convexity of KL divergence in its first argument, $D_{\mathrm{KL}}\left(q_{z|x}\,\|\,p_{z|x}\right)$ is strictly concave in $q_{z\mid x}$.
Focusing on the second KL term, from Equation 29 we have that
$D_{\mathrm{KL}}\left(q_{z|x}q_{w|x,y}\,\|\,p_{z,w|x,y}\right)=H(q_{z\mid x})+H(q_{w\mid x,y})+\mathbb{E}_{q_{z|x}}\mathbb{E}_{q_{w|x,y}}\log p(z,w\mid x,y),$ (32)
where $H(q_{z\mid x})$ is strictly concave in $q_{z|x}$.
Recall that we assumed that $p(w|y)$ depends on $q(z|x)$. Under our model
$p(x,y,z,w)=p(y)p(w\mid y)p(z)p(x\mid z,w),$ (33)
yielding
$p\left(z,w,\mid x,y\right)=p(w\mid y)\frac{p(z)p(x\mid z,w)}{p(x\mid y)}.$ (34)
Hence,
$\mathbb{E}_{q_{z|x}}\mathbb{E}_{q_{w|x,y}}\log p(z,w\mid x,y)=\mathbb{E}_{q_{z|x}}\mathbb{E}_{q_{w|x,y}}\left[\log p(w\mid y)+\log\frac{p(z)p(x\mid z,w)}{p(x\mid y)}\right],$
where $p(w\mid y)=\mathcal{N}\left(w;\mu_{y},I\right)$ with $\mu_{y}=\mathbb{E}_{p_{x|y}}\left[\mathbb{E}_{q_{z|x}}[z]\right]$. Therefore,
$\mathbb{E}_{q_{z|x}}\mathbb{E}_{q_{w|x,y}}\left[\log p(w\mid y)\right]=-\frac{1}{2}\left[d\log(2\pi)+\mathbb{E}_{q_{z|x}}\mathbb{E}_{q_{w|x,y}}\left\|w-\mu_{y}\right\|^{2}\right]$ (35)
where $\left\|w-\mu_{y}\right\|^{2}$ is a quadratic form in $\mu_{y}$, which is linear in $q_{z|x}$, and thus $\mathbb{E}_{q_{z|x}}\mathbb{E}_{q_{w|x,y}}\left[\log p(w\mid y)\right]$ is strictly concave in $q_{z|x}$. Hence, $D_{\mathrm{KL}}\left(q_{z|x}q_{w|x,y}\,\|\,p_{z,w|x,y}\right)$ is strictly concave in $q_{z|x}$, and thus so is $\mathbb{E}_{p_{x,y}}[\ell_{z}]$. ∎
Appendix C Supplementary Figures
[…]
Supplementary Figure 1: Encoder-decoder structures for previous approaches. (a) CSVAE. (b) DIVA - CCVAE.
Supplementary Figure 2: Comparison of approximate variational posteriors against the true posterior for latent variables $z,w$ for different values of $x$ with $y=0$ (top) and $y=1$ (bottom).
Supplementary Figure 3: Colored MNIST results for no noise.
Supplementary Figure 4: Colored MNIST visual results across the remaining noise levels.
Supplementary Figure 5: Reconstruction performance for other models on the CelebA-Glasses dataset. Top: Original samples from the data. Bottom: Reconstructions by the given model.
Supplementary Figure 6: Embeddings obtained by other models on the Kang dataset. For each block, top (resp. bottom) rows are $z$ (resp. $w$) embeddings, while left (resp. right) columns are colored by cell type (resp. stimulation).
Appendix D Supplementary tables for experimental results
Supplementary Table 1: RMSE for the Colored MNIST dataset without any label noise.
Marginal RMSE ($p=0$) $\downarrow$
CSVAE - No Adv. $\mathbf{0.064\pm 0.002}$
CSVAE $0.079\pm 0.008$
HCSVAE - No Adv. $0.094\pm 0.004$
HCSVAE $0.079\pm 0.030$
DIVA $0.065\pm 0.005$
CCVAE $0.065\pm 0.006$
DisCoVR (Ours) $\mathbf{0.064\pm 0.000}$
Supplementary Table 2: RMSE calculated between the estimated and true marginal across different levels of label noise on the Colored MNIST dataset. $p$ defines label flip probability. Bold denotes best performance.
Marginal RMSE $\downarrow$
$p=0.1$ $p=0.2$ $p=0.3$ $p=0.4$
CSVAE - No Adv. $0.141\pm 0.002$ $0.141\pm 0.003$ $0.142\pm 0.002$ $0.143\pm 0.002$
CSVAE $0.135\pm 0.022$ $0.152\pm 0.018$ $0.181\pm 0.007$ $0.173\pm 0.008$
HCSVAE - No Adv. $0.150\pm 0.001$ $0.150\pm 0.000$ $0.151\pm 0.000$ $0.151\pm 0.001$
HCSVAE $0.139\pm 0.003$ $0.141\pm 0.001$ $0.141\pm 0.001$ $0.141\pm 0.001$
DIVA $0.115\pm 0.011$ $0.102\pm 0.013$ $0.106\pm 0.010$ $0.113\pm 0.014$
CCVAE $0.092\pm 0.002$ $103\pm 0.014 0.0$$99\pm 0.011 0.0$$092\pm 0.005$
DisCoVR (Ours) $\mathbf{0.073\pm 0.001}$ $\mathbf{0.083\pm 0.004}$ $\mathbf{0.087\pm 0.002}$ $\mathbf{0.087\pm 0.001}$
Supplementary Table 3: Model performances on the CelebA-Glasses dataset. Bold denotes best performance.
$I(z;w)\downarrow$ NLL ($\downarrow$)
CSVAE - No Adv. $0.048\pm 0.014$ $137.522\pm 0.155$
CSVAE $0.079\pm 0.029$ $145.989\pm 0.336$
HCSVAE - No Adv. $0.055\pm 0.012$ $131.813\pm 0.21$
HCSVAE $0.055\pm 0.014$ $137.319\pm 0.265$
DIVA $0.188\pm 0.028$ $143.528\pm 0.02$
CCVAE $0.083\pm 0.022$ $\mathbf{131.764\pm 0.006}$
DisCoVR (Ours) $\mathbf{0.030\pm 0.011}$ $135.677\pm 0.007$
DisCoVR - Common (Ours) $374.114\pm 0.05$
Appendix E Additional disentanglement metrics
We provide an extended disentanglement assessment using multiple metrics. Because mutual information is difficult to estimate reliably, we report two estimators—MINE and kNN. Although their absolute values differ, the relative rankings of the methods remain consistent as can be seen in the ranking tables. In addition to these mutual-information estimates, we also report the following metrics, which quantifying the level of label information captured by $w$ compared to $z$ :
Mutual Information Gap (MIG)
$\operatorname{MIG}(w;z)=\frac{I(y;w)-I(y;z)}{H(y)}$
Mutual Information Completeness (MIC)
$\operatorname{MIC}(w;z)=\frac{I(y;w)}{I(y;w)+I(y;z)}$
E.1 Parametric model
CSVAE and its variants impose a fully separable prior, thereby forcing separability even when the true latent structure is not separable (see Table 1). In contrast, DisCoVR learns informative conditional embeddings that closely track the true posterior without requiring ground-truth knowledge of a truncated or fully separable prior, and it outperforms both DIVA and CCVAE.
Replacing the prior in DisCoVR with a fully separable predefined prior on $w$ yields consistent embeddings with the ground-truth structure while retaining the benefits of separability.
Supplementary Table 4: Additional disentanglement metrics calculated with kNN mutual information estimation for the parametric model dataset with $k=20$. Bold indicates closest to true posterior within group.
Assumption Model $I(y;z)$ $I(y;w)$ $I(w;z)$ $\operatorname{MIG}(w;z)$ $\operatorname{MIC}(w;z)$ $I(w;z\mid y)$
[…]
HCSVAE - N.A. $0.000\pm 0.000$ $\mathbf{0.643\pm 0.000}$ $000\pm 0.001 0.0$$072\pm 0.000$ $1.000\pm 0.000$ $0.000\pm 0.000$
HCSVAE $0.000\pm 0.000$ $\mathbf{0.643\pm 0.001}$ $001\pm 0.001 0.0$$072\pm 0.000$ $1.000\pm 0.000$ $0.000\pm 0.001$
DisCoVR (CSVAE prior) $0.000\pm 0.000$ $\mathbf{0.643\pm 0.000}$ $051\pm 0.007 0.0$$072\pm 0.000$ $1.000\pm 0.000$ $\mathbf{0.031\pm 0.005}$
Flexible DIVA $021\pm 0.042 0.0$$91\pm 0.046 0.0$$00\pm 0.000 0.0$$008\pm 0.010$ $800\pm 0.400 0.0$$000\pm 0.000$
CCVAE $\mathbf{0.022\pm 0.043}$ $90\pm 0.045 0.000$$0\pm 0.000 0.0$$008\pm 0.010$ $800\pm 0.400 0.0$$000\pm 0.000$
DisCoVR (our prior) $0.010\pm 0.006$ $\mathbf{0.151\pm 0.007}$ $\mathbf{0.108\pm 0.029}$ $\mathbf{0.016\pm 0.001}$ $\mathbf{0.938\pm 0.035}$ $\mathbf{0.072\pm 0.020}$
Fully Separable Posterior (no truncation) $057\pm 0.001 0.0$$057\pm 0.000$ $144\pm 0.003 0.0$$000\pm 0.000$ $499\pm 0.006 0.0$$090\pm 0.002$
True Posterior $0.058\pm 0.003$ $0.643\pm 0.000$ $144\pm 0.005 0.0$$066\pm 0.000$ $917\pm 0.004 0.0$$055\pm 0.003$
Supplementary Table 5: Additional disentanglement metrics calculated with MINE mutual information estimation for the parametric model dataset. Bold indicates closest to true posterior within group.
Assumption Model $I(y;z)$ $I(y;w)$ $I(w;z)$ $\operatorname{MIG}(w;z)$ $\operatorname{MIC}(w;z)$ $I(w;z\mid y)$
[…]
CSVAE $0.033\pm 0.055$ $\mathbf{0.526\pm 0.045}$ $031\pm 0.045 0.0$$055\pm 0.010$ $945\pm 0.091 0.0$$009\pm 0.005$
HCSVAE - N.A. $0.000\pm 0.000$ $543\pm 0.018 0.0$$00\pm 0.000 0.0$$061\pm 0.002$ $1.000\pm 0.000$ $0.000\pm 0.000$
HCSVAE $0.000\pm 0.000$ $543\pm 0.022 0.0$$00\pm 0.000 0.0$$061\pm 0.002$ $1.000\pm 0.000$ $0.001\pm 0.000$
[…]
Flexible DIVA $0.027\pm 0.053$ $113\pm 0.056 0.0$$01\pm 0.001 0.0$$010\pm 0.012$ $798\pm 0.398 0.0$$001\pm 0.000$
CCVAE $0.026\pm 0.052$ $115\pm 0.058 0.0$$01\pm 0.001 0.0$$010\pm 0.012$ $799\pm 0.399 0.0$$001\pm 0.000$
DisCoVR (ours) $\mathbf{0.037\pm 0.006}$ $\mathbf{0.176\pm 0.008}$ $\mathbf{0.109\pm 0.025}$ $\mathbf{0.016\pm 0.001}$ $\mathbf{0.825\pm 0.026}$ $\mathbf{0.073\pm 0.018}$
Fully Separable Posterior (no truncation) $084\pm 0.003 0.0$$083\pm 0.003$ $137\pm 0.004 0.0$$000\pm 0.000$ $497\pm 0.009 0.0$$088\pm 0.003$
True Posterior $0.085\pm 0.005$ $0.493\pm 0.011$ $139\pm 0.005 0.0$$046\pm 0.002$ $853\pm 0.009 0.0$$057\pm 0.004$
Supplementary Table 6: Rank (1 = closest to True Posterior) of each method with respect to the true posterior for metrics calculated with kNN mutual information estimation with $k=20$. Colors indicate rank within each block: red = worse (farther), green = better (closer).
Assumption Model $I(y;z)$ $I(y;w)$ $I(w;z)$ $\operatorname{MIG}(w;z)$ $\operatorname{MIC}(w;z)$ $I(w;z\mid y)$
[…]
Supplementary Table 7: Rank (1 = closest to True Posterior) of each method with respect to the True Posterior for metrics calculated with MINE mutual information estimation. Colors indicate rank within each block: red = worse (farther), green = better (closer).
Assumption Model $I(y;z)$ $I(y;w)$ $I(w;z)$ $\operatorname{MIG}(w;z)$ $\operatorname{MIC}(w;z)$ $I(w;z\mid y)$
[…]
Fully Separable DisCoVR (CSVAE prior) $cellcolorUpper!55!Lower 3 \cellcolorUpper$!70!Lower 2 $cellcolorUpper!55!Lower 3 \cellcolorUpper$!85!Lower 1
[…]
E.2 Noisy Swiss Roll
When the observed labels are noisy, DisCoVR outperforms other methods, obtaining embeddings close to the ground truth.
Supplementary Table 8: Additional disentanglement metrics calculated with kNN mutual information estimation for the Noisy Swiss Roll ($p=0.3$) dataset with $k=20$. Bold indicates closest to ground truth within group.
Assumption Model $I(y;z)$ $I(y;w)$ $I(w;z)$ $\operatorname{MIG}(w;z)$ $\operatorname{MIC}(w;z)$ $I(w;z\mid y)$
Fully Separable CSVAE - N.A. $0.041\pm 0.007$ $0.525\pm 0.221$ $362\pm 0.180 0.0$$057\pm 0.026$ $0.888\pm 0.098$ $0.266\pm 0.152$
[…]
HCSVAE - N.A. $0.029\pm 0.007$ $642\pm 0.000 0.0$$65\pm 0.013 0.0$$072\pm 0.001$ $957\pm 0.010 0.0$$009\pm 0.019$
[…]
Flexible DIVA $034\pm 0.013 0.0$$036\pm 0.011$ $2.633\pm 0.360$ $0.000\pm 0.003$ $0.515\pm 0.159$ $2.185\pm 0.332$
CCVAE $040\pm 0.015 0.0$$030\pm 0.007$ $2.952\pm 0.124$ $0.001\pm 0.003$ $0.447\pm 0.153$ $2.462\pm 0.118$
DisCoVR (ours) $\mathbf{0.000\pm 0.000}$ $\mathbf{0.049\pm 0.002}$ $\mathbf{0.029\pm 0.011}$ $\mathbf{0.006\pm 0.000}$ $\mathbf{1.000\pm 0.000}$ $\mathbf{0.014\pm 0.008}$
Noisy Ground Truth $000\pm 0.000 0.0$$55\pm 0.002 0.0$$00\pm 0.000 0.0$$007\pm 0.000$ $1.000\pm 0.000$ $0.000\pm 0.000$
Supplementary Table 9: Additional disentanglement metrics calculated with MINE mutual information estimation for the Noisy Swiss Roll ($p=0.3$) dataset. Bold indicates closest to ground truth within group.
Assumption Model $I(y;z)$ $I(y;w)$ $I(w;z)$ $\operatorname{MIG}(w;z)$ $\operatorname{MIC}(w;z)$ $I(w;z\mid y)$
CSVAE - N.A. $0.046\pm 0.022$ $422\pm 0.186 0.0$$50\pm 0.020 0.0$$044\pm 0.023$ $834\pm 0.184 0.0$$029\pm 0.017$
[…]
HCSVAE - N.A. $0.023\pm 0.014$ $585\pm 0.011 0.0$$26\pm 0.012 0.0$$066\pm 0.002$ $963\pm 0.021 0.0$$006\pm 0.002$
[…]
Noisy Ground Truth $000\pm 0.000 0.0$$24\pm 0.018 0.0$$00\pm 0.001 0.0$$003\pm 0.002$ $985\pm 0.048 0.0$$002\pm 0.001$
Supplementary Table 10: Rank (1 = closest to Ground Truth) of each method with respect to the Ground Truth for metrics calculated with kNN mutual information estimation with $k=20$ on the Noisy Swiss Roll ($p=0.3$) dataset. Colors indicate rank within each block: red = worse (farther), green = better (closer).
Assumption Method $I(y;z)$ $I(y;w)$ $I(w;z)$ $\operatorname{MIG}(w;z)$ $\operatorname{MIC}(w;z)$ $I(w;z\mid y)$
[…]
Supplementary Table 11: Rank (1 = closest to Ground Truth) of each method with respect to the Ground Truth for metrics calculated with MINE mutual information estimation on the Noisy Swiss Roll ($p=0.3$) dataset. Colors indicate rank within each block: red = worse (farther), green = better (closer).
Assumption Model $I(y;z)$ $I(y;w)$ $I(w;z)$ $\operatorname{MIG}(w;z)$ $\operatorname{MIC}(w;z)$ $I(w;z\mid y)$
[…]
Appendix F Ablations on model components
We evaluate the contribution of each model component by examining two variations: (1) applying the classifier directly to $z$, and (2) replacing the conditional prior on $w$ with a standard Gaussian.
When training the classifier directly on $z$ we were able achieve results qualitatively similar to those obtained using the reconstruction $\hat{x}$, but doing so requires substantially more parameter tuning.
An unconditional standard Gaussian prior for $w$, causes $w$ to collapse into a representation redundant with $z$, removing meaningful separation.
Supplementary Figure 7: Ablation study on the Noisy Swiss Roll ($p=0.3$) dataset.
Supplementary Figure 8: Ablation study on the Noisy Colored MNIST ($p=0.3$) dataset. For each setting: left column denotes coloring by noisy labels, right column denotes coloring by digit (shape, not included in the label).
Appendix G Additional experiment on CelebA-Hats
We performed an additional experiment on the CelebA dataset, with the attribute Wearing_hatWearing\_hat denoting the $y$ label. Supplementary Table 12 outlines the results of this experiment. DisCoVR is the only method that exhibits high disentanglement for $z,w$ without compromising reconstruction quality.
Supplementary Table 12: Model performances of a single experiment on CelebA-Hats. Bold denotes best performance.
$I(z;w)\downarrow$ NLL ($\downarrow$)
CSVAE - N.A. 0.360 653.537
CSVAE 0.213 351.082
HCSVAE - N.A. 0.135 2608.442
HCSVAE 0.192 673.674
DIVA 0.553 356.090
CCVAE 0.856 347.940
DisCoVR (Ours) 0.059 353.271
DisCoVR (Ours) - Common - 437.144
Appendix H Implementation details
H.1 Considerations and reproducibility
We run all experiments on a single H100 GPU. Reported means and standard deviations for tables are conducted over 10 repetitions of the experiment with different random seeds. All models are trained using the AdamW (loshchilov2018decoupled) optimizer until validation loss stops decreasing for 50 epochs. Wherever provided, we use mutual information neural estimation (MINE, pmlr-v80-belghazi18a) and k-Nearest Neighbor (kNN) mutual information estimation Kraskov2011 to obtain mutual information estimates. For Naive Bayes classifiers, we use the implementation provided by scikit-learn (scikit-learn). To use ideal hyperparameters for each method, we consult the original implementation whenever possible, and conduct a simple grid-search to produce originally described model behavior. Implementations of all methods compared in this study, including DisCoVR, as well as code to reproduce our results, is attached to this submission and will be made public upon acceptance. Models compared in the study admit a weighting term for each term in the loss function, of which most are shared across different approaches. We use the following shorthands for each of the terms:
$\text{Rec.}\rightarrow\mathbb{E}_{q_{z|x}}\left[\mathbb{E}_{q_{w|x,y}}[\log p(x\mid z,w)]\right]$
$D_{\mathrm{KL}}(Z)\rightarrow D_{\mathrm{KL}}(q_{z|x}\,\|\,p_{z})$
$D_{\mathrm{KL}}(W)\rightarrow D_{\mathrm{KL}}\left(q_{w|x,y}\,\|\,p_{w\mid y}\right)$
$\text{Adv.}\rightarrow-\mathbb{E}_{q_{z|x}}[\log g(y\mid z)]$
$\text{Class.}\rightarrow\mathbb{E}_{q_{w|x,y}}[\log q(y\mid w)]$
$\text{Rec. - ($Z$)}\rightarrow\mathbb{E}_{q_{z|x}}[\log p(x\mid z)]$
Below, we provide additional details for the hyperparameters used in each experiment, and any other external resources used to obtain the corresponding sections’ results. In addition, we include details regarding runtime and memory footprint of running experiments with the models included in our study.
Supplementary Table 13: Time spent per epoch during training for each dataset.
P.M. N.S.R CMNIST CelebA scRNA-seq
CSVAE - N.A. 10.91s 8.02s 14s 54.43s 4.78s
CSVAE 12.71s 8.98s 14.41s 74.68s 4.99s
HCSVAE - N.A. 15.6s 10.92s 17s 44.3s 6.34s
HCSVAE 16.51s 11.8s 16.8s 68.53s 5.3s
DIVA 11.1s 8s 18.59s 48.96s 3.55s
CCVAE 12.1s 8.44s 17.9s 49.65s 5.25s
DisCoVR(Ours) 12.18s 8.73s 21.8s 109.86s 6.09s
Supplementary Table 14: Model inference time for a single batch for each dataset.
P.M. N.S.R CMNIST CelebA scRNA-seq
CSVAE - N.A. 72ms 28ms 57ms 329ms 100ms
CSVAE 40ms 29ms 59ms 187ms 96ms
HCSVAE - N.A. 36ms 30ms 55ms 214ms 95ms
HCSVAE 37ms 39ms 52ms 191ms 119ms
DIVA 25ms 26ms 60ms 154ms 42ms
CCVAE 27ms 28ms 42ms 163ms 93ms
DisCoVR(Ours) 32ms 27ms 63ms 205ms 123ms
Supplementary Table 15: Memory footprint of running an experiment for each dataset.
P.M. N.S.R CMNIST CelebA scRNA-seq
CSVAE - N.A. 53 MiB 255 MiB 1988 MiB 4868 MiB 298 MiB
CSVAE 253 MiB 255 MiB 2378 MiB 4812 MiB 300 MiB
HCSVAE - N.A. 254 MiB 255 MiB 2558 MiB 4588 MiB 292 MiB
HCSVAE 253 MiB 256 MiB 2998 MiB 4466 MiB 294 MiB
DIVA 253 MiB 255 MiB 2634 MiB 4996 MiB 300 MiB
CCVAE 253 MiB 255 MiB 3066 MiB 4998 MiB 300MiB
DisCoVR (Ours) 254 MiB 257 MiB 3612 MiB 7078 MiB 308MiB
H.1.1 Parametric model
for the parametric model, we consider $z,w\in\mathbb{R}$ and use multi-layer perceptrons (MLPs) with $n_{hidden}=2,d_{hidden}=8$ to parameterize approximate posteriors, the generative model and classifiers. For all models, we use learning rate $\gamma=0.001$. A more detailed table of model-specific loss weights is provided in Supplementary Table 16.
Supplementary Table 16: Loss weights for the parametric model experiment.
Rec. $D_{\mathrm{KL}}(Z)$ $D_{\mathrm{KL}}(W)$ Adv. Class. Rec. - ($Z$)
CSVAE - No Adv. 1 1 1
CSVAE 2.5 1 0.5 20
HCSVAE - No Adv. 1 1 0.5
HCSVAE 2.5 1 0.5 20
DIVA 1 1 1 1
CCVAE 1 1 1 1
DisCoVR (Ours) 0.75 0.9 0.2 0.8 0.25
Supplementary Table 17: K-Means NMI for embeddings across stimulation ($y$) and cell type (common structure).
$w$ - Stimulation ($\uparrow$) $z$ - Cell Type ($\uparrow$) $z$ - Stimulation ($\downarrow$)
CSVAE - No Adv. $0.949\pm 0.003$ $0.702\pm 0.015$ $0.187\pm 0.0$
[…]
HCSVAE - No Adv. $0.933\pm 0.006$ $628\pm 0.016 0.0$$091\pm 0.002$
HCSVAE $0.931\pm 0.005$ $433\pm 0.001 0.0$$003\pm 0.0$
DIVA $0.801\pm 0.0$ $628\pm 0.011 0.0$$056\pm 0.0$
CCVAE $0.604\pm 0.0$ $0.683\pm 0.016$ $0.103\pm 0.0$
DisCoVR (Ours) $\mathbf{0.946\pm 0.002}$ $\mathbf{0.688\pm 0.0}$ $\mathbf{0.002\pm 0.0}$
H.1.2 Noisy Swiss roll
For this experiment, we consider $z,w\in\mathbb{R}^{2}$ and use MLPs with $n_{hidden}=2,d_{hidden}=128$ to parameterize approximate posteriors, the generative model and classifiers. For all models, we use learning rate $\gamma=0.001$. A more detailed table of model-specific hyperparameters is provided in Supplementary Table 18.
Supplementary Table 18: Loss weights for the noisy Swiss roll experiment.
Rec. $D_{\mathrm{KL}}(Z)$ $D_{\mathrm{KL}}(W)$ Adv. Class. Rec. - ($Z$)
CSVAE - No Adv. 20 0.2 1
CSVAE 20 0.2 1 50
HCSVAE - No Adv. 20 0.2 1
HCSVAE 20 0.5 1 50
DIVA 20 0.2 0.2 1
CCVAE 20 0.2 0.2 1
DisCoVR (Ours) 0.9 0.2 0.2 8 0.1
H.1.3 Noisy colored MNIST
For this experiment, we consider $z\in\mathbb{R}^{20}$, $w\in\mathbb{R}^{2}$ and use convolutional neural networks (CNNs) to parameterize approximate posteriors and the generative model. For this example, DisCoVR can support $z,w$ with different sizes, by parameterizing $p(w\mid y)$ through neural networks. For all models, we use learning rate $\gamma=0.0001$. We detail the architectures and model-specific hyperparameters in Supplementary Tables 19 - 22. All other neural networks are formulated as MLPs with $n_{hidden}=2,d_{hidden}=4096$.
Supplementary Table 19: Image encoder architecture for noisy colored MNIST. Parameters for Conv2d are input / output channels. Parameters for MaxPool2D are kernel size and stride. Parameter for the linear layer is the output size. For variances, outputs are passed through an additional Softplus layer to ensure non-negativity.
Block Details
1 Conv2d(3,32) + BatchNorm2D + ReLU
2 Conv2d(32,32) + BatchNorm2D + ReLU + MaxPool2D(2,2)
3 Conv2d(32,64) + BatchNorm2D + ReLU + MaxPool2D(2,2)
4 Conv2d(64,128) + BatchNorm2D + ReLU + MaxPool2D(2,2)
5 Linear(4096) + BatchNorm1D + ReLu
6 Linear(4096) + BatchNorm1D + ReLu
7 Linear($d_{latent}$)
Supplementary Table 20: Image decoder architecture for noisy colored MNIST. Parameters for Conv2d are input / output channels. Parameters for MaxPool2D are kernel size and stride. Parameter for the linear layer is the output size.
Block Details
1 Linear(4096) + BatchNorm1D + ReLu
2 Linear(4096) + BatchNorm1D + ReLu
3 Linear(1152) + Unflatten
4 Upsample(2) + Conv2d(128, 64) + BatchNorm2D + ReLU
5 Upsample(2) + Conv2d(64, 32) + BatchNorm2D + ReLU
6 Upsample(2) + Conv2d(32, 32) + BatchNorm2D + ReLU
7 Conv2d(32, 3) + Sigmoid
Supplementary Table 21: Latent classifier architecture for noisy colored MNIST. Outputs parameterize logits of class probabilities.
Block Details
1 Linear(4096) + BatchNorm1D + ReLu
2 Linear(4096) + BatchNorm1D + ReLu
3 Linear(2)
Supplementary Table 22: Loss weights for the noisy colored MNIST experiment.
Rec. $D_{\mathrm{KL}}(Z)$ $D_{\mathrm{KL}}(W)$ Adv. Class. Rec. - ($Z$)
CSVAE - No Adv. 1 0.0001 1
CSVAE 1 0.0001 1 1
HCSVAE - No Adv. 1000 0.0001 1
HCSVAE 10000 0.0001 1 1
DIVA 1 0.0001 0.0001 1
CCVAE 1 0.0001 0.0001 1
DisCoVR (Ours) 0.5 0.0001 0.0001 0.1 0.5
H.1.4 CelebA-glasses
Motivated by the previous application of CSVAE, our choices follow those outlined in pmlr-v48-larsen16. We provide a detailed table of model-specific hyperparameters in Supplementary Table 23:
Supplementary Table 23: Loss weights for the CelebA-Glasses experiment.
Rec. $D_{\mathrm{KL}}(Z)$ $D_{\mathrm{KL}}(W)$ Adv. Class. Rec. - ($Z$)
CSVAE - No Adv. 1 0.0001 1
CSVAE 1000 0.0001 1 1
HCSVAE - No Adv. 1000 0.0001 1
HCSVAE 10000 0.0001 1 1
DIVA 100000 0.0001 0.0001 1
CCVAE 100000 0.0001 0.0001 1
DisCoVR (Ours) 1000000 0.0001 0.0001 2000 100000
H.1.5 scRNA-Seq
Following on the previous applications by scVI, we use $z\in\mathbb{R}^{10}$, $w\in\mathbb{R}^{2}$. For DisCoVR  we match the dimensions and use $w\in\mathbb{R}^{10}$ to avoid parameterizing the prior $p(w\mid z,y)$ with an additional neural network. We use MLPs with $n_{hidden}=1,d_{hidden}=128$ to parameterize approximate posteriors, the generative model and classifiers. We calculate K-Means NMI through scikit-learn (scikit-learn) by calling the normalized_mutual_info_score function with the original labels and the clusterings obtained by running KMeans on (1) the entire latent embedding and (2) single dimensions of the embedding and report the highest score. A more detailed table of model-specific hyperparameters is provided in Supplementary Table 24:
Supplementary Table 24: Loss weights for the scRNA-seq experiment.
Rec. $D_{\mathrm{KL}}(Z)$ $D_{\mathrm{KL}}(W)$ Adv. Class. Rec. - ($Z$)
CSVAE - No Adv. 1 0.0001 1
CSVAE 1 0.0001 1 100
HCSVAE - No Adv. 1 0.0001 1
HCSVAE 1 0.0001 1 100
DIVA 1 0.0001 0.0001 1
CCVAE 1 0.0001 0.0001 1
DisCoVR (Ours) 0.9 0.0001 0.0001 100 0.1
H.2 Summary of the scVI generative model for 5.2.3
Given batch key $b$ and $G$ genes, the generative model of scVI for a single cell $x_{i}\in\mathbb{N}^{G}$ is formulated as:
$z_{i}\sim\mathcal{N}(0,1)$
$\rho_{i}=f_{\theta}(z_{i},b_{i})$
$\pi_{ig}=h_{\phi}^{g}(z_{i},b_{i})$
$x_{ig}\sim\text{ZINB}(l_{i}\rho_{i},\theta_{g},\pi_{ig})$
Here, $g$ indexes genes, $l_{i}=\sum_{g}x_{ig}$ denotes the total number of counts for a single cell, $z_{i}$ denotes the latent representation of the cell, and $\rho_{i}$ denotes the normalized expression of the cell. $f_{\theta}$ is formulated as a neural network with a final softmax layer. $h_{\phi}$ is a neural network used to parameterize zero-inflation probabilities for the generative zero-inflated negative binomial (ZINB) distribution. As such, for a single batch, the formulation of scVI is equivalent to the VAE with a ZINB likelihood. While all other models can be extended easily, DisCoVR requires reconstructions as a proxy for the adversarial loss. For this formulation, we directly treat the normalized expressions $\rho_{i}$ as the adversarial reconstructions $\hat{x}$.